### v2

# Reverse Diffusion Sequential Monte Carlo Samplers

Abstract
We propose a novel sequential Monte Carlo (SMC) method for sampling from unnormalized target distributions based on a reverse denoising diffusion process. While recent diffusion-based samplers simulate the reverse diffusion using approximate score functions, they can suffer from accumulating errors due to time discretization and imperfect score estimation. In this work, we introduce a principled SMC framework that formalizes diffusion-based samplers as proposals while systematically correcting for their biases. The core idea is to construct informative intermediate target distributions that progressively steer the sampling trajectory toward the final target distribution. Although ideal intermediate targets are intractable, we develop exact approximations using quantities from the score estimation-based proposal, without requiring additional model training or inference overhead. The resulting sampler, termed Reverse Diffusion Sequential Monte Carlo, enables consistent sampling and unbiased estimation of the target’s normalization constant under mild conditions. We demonstrate the effectiveness of our method on a range of synthetic targets and real-world Bayesian inference problems.  [^1] Our code is available at https://github.com/LuhuanWu/RDSMC.
1 Introduction
Sampling from unnormalized target distributions is a fundamental problem in many applications, ranging from Bayesian inference [andrieu2003introduction, robert2004metropolis] to simulating molecular systems [frenkel2000molecular]. Classical methods like Markov chain Monte Carlo (MCMC) simulate a Markov chain with the target as its stationary distribution, but they can suffer from slow mixing and difficulty traversing between modes for complex distributions. Particle methods such as importance sampling generate exact samples in a large compute limit; yet they struggle with the curse of dimensionality [chatterjee2018sample]. Alternatively, variational inference (VI) [blei2017variational] casts inference as an optimization task, though its success depends on the expressiveness of the variational family and the complexity of the optimization landscape.
Recently, diffusion models have emerged as a powerful approach for sampling from complex distributions [ho2020denoising, song2020score]. They define a forward noising process that gradually transforms a complex target distribution into a simple base distribution. A reverse denoising process then reconstructs target samples by simulating the dynamics backward in time, starting from the base distribution and guided by a time-dependent score function. In the generative modeling setting, this score function is approximated by a neural network trained on samples from the target distribution. However, in the sampling context, such training data are unavailable and only an unnormalized target density is accessible.
Recent works on diffusion-based samplers explore alternative ways to approximate the score function directly from the target density, enabling sampling without access to training data. One line of works, known as diffusion Monte Carlo (MC) samplers, estimates the score function using MC methods. huang2023reverse, grenioux2024stochastic consider Langevin-style MC algorithms, but they rely on a good initialization of the reverse diffusion process to ensure theoretical guarantees. he2024zeroth propose an alternative scheme based on rejection sampling which relaxes prior assumptions and improves sampling efficiency in low-dimensional regimes. These approaches demonstrate both theoretical and empirical advantages over conventional MCMC methods, particularly for multi-modal distributions.
In contrast to relying on MC estimation during sampling time, a complementary line of research trains a neural network in advance to approximate the score function. To this end, some works propose new score matching objectives [phillips2024particle, akhound2024iterated, ouyang2024bnem], while variational approaches instead optimize divergences between forward and reverse diffusion processes [zhang2021path, berner2022optimal, richter2023improved, vargas2023denoising, vargas2023transport].
While promising, diffusion-based samplers suffer from two sources of bias: discretization error in simulating the reverse diffusion process and approximation error in the estimated or learned score function. An exception is phillips2024particle which mitigates the bias of a trained diffusion-based sampler using Sequential Monte Carlo (SMC), a general inference tool for sequential models [naesseth2019elements, chopin2020introduction]. However, such training-based methods remain computationally complex compared to classical sampling methods and often rely on special neural network preconditioning [he2025no].
To address these challenges, we develop a new diffusion-based sampler, Reverse Diffusion Sequential Monte Carlo (RDSMC) for sampling from unnormalized target distributions, which is training-free and admits theoretical guarantees. Inspired by prior work, we formalize diffusion MC samplers as proposal mechanisms within an SMC framework. At a high level, RDSMC generates multiple particles from the reverse diffusion dynamics using MC-based score estimates. To correct the bias in proposals, we introduce intermediate target distributions that guide resampling of particles at each step, progressively steering them toward the final target distribution of interest.
Crucially, our intermediate targets are efficiently computed using byproducts of MC-based score estimates, incurring no additional cost. Moreover, they form an exact approximation to the ideal intermediate targets that maximize sampling efficiency, defined by the marginal distributions of an extended final target [naesseth2019elements, andrieu2010particle]. This design helps particles stay closely aligned with the final target throughout the sampling process. In contrast, phillips2024particle learn neural network-based surrogates that require additional training and attain the ideal target only at the final step.
RDSMC belongs to a class of nested SMC methods [naesseth2015nested, naesseth2019high], inheriting the standard SMC guarantees. In particular, it produces asymptotically exact samples from the target in the limit of many particles, and provides an unbiased estimate of the normalization constant for any fixed size of particles.
Our contributions are summarized as follows:
  • We propose a new SMC algorithm, Reverse Diffusion Sequential Monte Carlo (RDSMC), based on the reverse diffusion process for sampling from unnormalized distributions.
  • RDSMC is training-free and extends existing diffusion MC samplers to achieve asymptotically exact sampling and provide unbiased estimates of the normalization constant, while incurring almost no computational overhead given the same number of final samples.
  • Empirically, RDSMC outperforms or matches existing diffusion MC samplers and classical geometric annealing-based Annealed Importance Sampling (AIS) [neal2001annealed] and SMC samplers [del2006sequential] on synthetic targets and Bayesian logistic regression benchmarks.
2 Background
Diffusion models.
Diffusion models [ho2020denoising, song2020score] evolve a complex target distribution $\pi(x)$ into a simple base distribution $\pi_{1}(x)$, e.g. $\pi_{1}(x)=\mathcal{N}(0,1)$, via a forward stochastic differential equation (SDE)
$dx_{t}=f(t)x_{t}dt+g(t)dB_{t},\qquad t:0\to 1,$ (1)
where $f(t)$ and $g(t)$ are the drift and diffusion coefficients, and $B_{t}$ is the standard Brownian motion.
To generate samples from $\pi(x)$, we simulate a reverse SDE initialized from $x_{1}\sim\pi_{1}(x)$,
$dx_{t}$ $=\left[f(t)x_{t}-g(t)^{2}\nabla_{x_{t}}\log\pi_{t}(x_{t})\right]dt+g(t)d\bar{B}_{t},\qquad t:1\to 0,$ (2)
where $\nabla_{x_{t}}\log\pi_{t}(x_{t})$ is the score function of the marginal density $\pi_{t}(x_{t})$ at time $t$ induced by the forward process in Eq.˜1, and $\bar{B}_{t}$ is the reverse-time Brownian motion.
Diffusion MC samplers.
Building on the diffusion model paradigm, recent works use MC methods to simulate the reverse dynamics in Eq.˜2 whose terminal distribution corresponds to the target distribution $\pi\propto\tilde{\pi}$ of interest [huang2023reverse, grenioux2024stochastic, he2024zeroth]. While the score function $\nabla_{x_{t}}\log\pi_{t}(x_{t})$ is generally intractable, the key idea is to construct MC estimates via the denoising score identity [DSI, vincent2011connection, efron2011tweedie],
$\nabla_{x_{t}}\log\pi_{t}(x_{t})$ $=\int\frac{\alpha(t)x_{0}-x_{t}}{\sigma(t)^{2}}\pi(x_{0}\mid x_{t})dx_{0},$ (3)
where $\pi(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{0}\mid\alpha(t)x_{0},\sigma(t)^{2}\mathbb{I})$ is the denoising posterior, which treats the unnormalized target $\tilde{\pi}(x_{0})$ as the prior and the forward transition density $\mathcal{N}(x_{0}\mid\alpha(t)x_{0},\sigma(t)^{2}\mathbb{I})$ from Eq.˜1 as the likelihood. The coefficients $\alpha(t)$ and $\sigma(t)$ are determined by the drift and diffusion terms $f(t)$ and $g(t)$ (see §˜A.1 for details).
Eq.˜3 suggests that score estimation can be cast as a posterior inference problem. In practice, one can draw approximate samples from the denoising posterior to form an MC estimate of the score, which is then substituted into the reverse dynamics to generate samples from $\pi$. Other score identities beyond the DSI can also be leveraged (see §˜A.2).
Sequential Monte Carlo.
SMC [naesseth2019elements, chopin2020introduction] is a particle-based method for sampling from a sequence of distributions defined on variables $x_{0:T}$, terminating at a final target distribution of interest. We consider a reverse-time formulation, where SMC evolves a weighted collection of $N$ particles $\{x_{t}^{(i)},w_{t}^{(i)}\}_{i=1}^{N}$ from $t=T$ to 0, gradually approximating the final target.
An SMC sampler requires two key design choices [naesseth2019elements], a sequence of intermediate proposals, $q_{T}(x_{T})$ and $\{q_{t}(x_{t}\mid x_{t+1})\}_{t=0}^{T-1}$, and a sequence of (unnormalized) intermediate target distributions, $\{\gamma_{t}(x_{t:T})\}_{t=0}^{T}$, such that the final target $\gamma_{0}(x_{0:T})$ recovers the distribution of interest.
SMC initializes $N$ particles $x_{T}^{(i)}\sim q_{T}(x_{T})$ with weights $w_{T}^{(i)}\leftarrow\gamma_{T}(x_{T}^{(i)})/q_{T}(x_{T}^{(i)})$ for $i=1,\cdots,N$. Then for each step $t=T{-}1,\dots,0$ and particle $i=1,\dots,N$, it proceeds as follows:
  1. 1.
    resample ancestor $x_{t+1}^{(i)}\sim\text{Multinomial}\big(x_{t+1}^{(1:N)},w_{t+1}^{(1:N)}\big)$;
  2. 2.
    propagate particle $x_{t}^{(i)}\sim q_{t}\big(x_{t}\mid x_{t+1}^{(i)}\big)$;
  3. 3.
    compute weight $w_{t}^{(i)}\leftarrow\gamma_{t}(x_{t:T}^{(i)})\big/\left[\gamma_{+1}(x_{t+1:T}^{(i)})q_{t}(x_{t}^{(i)}\mid x_{t+1}^{(i)})\right]$.
The final set of weighted particles forms a discrete approximation to the final target $\gamma_{0}$, which is asymptotically exact given infinite particles under regularity conditions [chopin2020introduction]. However, the efficacy of SMC greatly depends on the choice of intermediate targets and proposals – the closer these intermediate distributions are to the marginals of the final target, the more effective the SMC sampler.
3 Method
Our goal is to sample from a target distribution $\pi(x)=\frac{1}{Z}\tilde{\pi}(x)$, where $\tilde{\pi}(x)$ is the unnormalized target and $Z=\int\tilde{\pi}(x)\mathrm{d}x$ is a generally intractable normalization constant. We develop Reverse Diffusion Sequential Monte Carlo (RDSMC), a diffusion-based SMC sampler targeting $\pi(x)$.
To enable sequential inference with SMC, we define an extended target distribution over a discretized trajectory $x_{0:T}=(x_{0},\dots,x_{T})$ of the forward process in Eq.˜1,
$\pi(x_{0:T})=\pi(x_{0})\prod_{t=1}^{T}\pi(x_{t}|x_{t-1}),$ (4)
where $0=\tau_{0}<\cdots<\tau_{T}=1$ are $T+1$ discretization times and $\pi(x_{t}|x_{t-1})$ is the forward transition density from time $\tau_{t-1}$ to $\tau_{t}$ induced by Eq.˜1 (see §˜A.1 for analytical expressions). Without loss of generality, we assume a uniform discretization step size $\delta=1/T=\tau_{t}-\tau_{t-1},\forall t=1,\cdots,T$.
This construction yields a sequence of intermediate targets $\{\pi(x_{t:T})\}_{t=T}^{0}$, whose final marginal $\pi(x_{0})$ recovers the desired target. These intermediate targets provide a natural setting for SMC, and in fact, are optimal intermediate gargets. However, they are intractable to sample from and to evaluate.
To address this challenge, we develop RDSMC leveraging the dual structure of diffusion processes, generating samples in the reverse direction while grounding their targets in the forward direction. In §˜3.1, we introduce a sequence of proposals based on the reverse diffusion process using MC score estimates. To correct the resulting proposal bias, in §˜3.2, we develop practical intermediate targets that form exact approximations to the optimal targets $\{\pi(x_{t:T})\}_{t=0}^{T}$ using byproducts of MC score estimates. Finally in §˜3.3, we present the full RDSMC algorithm and its theoretical guarantees.
Notation.
We slightly abuse notation by letting $x_{t}$ denote both the continuous-time variable $x_{t}$ for $t\in[0,1]$, as used in Eqs.˜1 and 2, and the discrete-time variable for $t\in\{0,\dots,T\}$, as used throughout this section. We denote $f_{t}\coloneqq f(\tau_{t}),g_{t}\coloneqq g(\tau_{t}),\alpha_{t}\coloneqq\alpha(\tau_{t}),\sigma_{t}\coloneqq\sigma(\tau_{t})$ for time-dependent diffusion coefficients from Eqs.˜1 and 3. For generality, we define $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} as the collection of all auxiliary random variables generated in the MC score estimation at step $t$.
3.1 Reverse diffusion proposal
We design an extended proposal distribution based on the reverse diffusion process with MC score estimates, jointly modeling diffusion dynamics $x_{t}$ and auxiliary randomness $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} from score estimation.
MC score estimation.
Given a sample $x_{t}$ at step $t$, we define a generic MC score estimator $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\approx\nabla_{x_{t}}\log\pi(x_{t})$ based on the DSI in Eq.˜3, where $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} is the estimation randomness associated with some sampling distribution $q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})$.
As an illustrative example, consider an importance sampling (IS) estimator (Algorithm˜2) targeting the posterior $\pi(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$. We draw $M$ importance samples $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}$} from the proposal $q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\mid x_{t})\coloneqq\mathcal{N}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\mid x_{t}/\alpha_{t},\sigma_{t}^{2}/\alpha_{t}^{2}\mathbb{I})$ for $m=1,\cdots,M$, and estimate the score by
$\nabla_{x_{t}}\log\pi(x_{t})$ $\approx s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\coloneqq\sum_{m=1}^{M}\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{M}w^{(m^{\prime})}}\cdot\frac{\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}-x_{t}}{\sigma_{t}^{2}},$ (5)
which corresponds to an MC approximation of the RHS of Eq.˜3. The importance weights $w^{(m)}$ are
$w^{(m)}=\frac{\tilde{\pi}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})\,\mathcal{N}(x_{t}\mid\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}},\sigma_{t}^{2}\mathbb{I})}{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\mid x_{t})},\qquad m=1,\cdots,M.$ (6)
In this case, the auxiliary randomness is to the collection of all importance samples ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}=\{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\}_{m=1}^{M}$.
In practice, we use a more sophisticated AIS scheme to improve the accuracy of score estimates (Algorithm˜3). More informative IS proposals or other MC methods such as SMC and rejection sampling may also be used; see §˜B.1 for further details.
Approximate reverse diffusion dynamics.
With the score estimates in place, we approximate the transition kernel of the reverse SDE in Eq.˜2 to define a conditional proposal for $x_{t}$,
$q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$ $\coloneqq\mathcal{N}(x_{t}\mid x_{t+1}-\left[f_{t+1}x_{t+1}-g_{t+1}^{2}s(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\right]\delta,g_{t+1}^{2}\delta).$ (7)
Extended proposal distributions.
The full sampling process starts by drawing $x_{T}$ from a tractable base distribution $q(x_{T})$, e.g. $q(x_{T})=\mathcal{N}(0,1)$. We then iteratively evolve $x_{t}$ for $t=T-1,\cdots,0$ using the reverse diffusion dynamics $q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$ in Eq.˜7 where auxiliary variable ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}\sim q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}\mid x_{t+1})$ is sampled in previous iteration to estimate the score $s(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$.
The overall sampling process defines a proposal over the extended space of $\{x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\}_{t=0}^{T}$,
$q(x_{0:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}})$ $\coloneqq q(x_{T})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})\prod_{t=0}^{T-1}q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t}).$ (8)
While sampling $\color[rgb]{0.3125,0.46875,0.62890625}u_{0}$} at the final step $t=0$ is not necessary, we retain it for notation simplicity.
This sampling procedure builds on prior work of huang2023reverse, grenioux2024stochastic, which estimate the score function using MCMC, and by he2024zeroth which use rejection sampling. We formalize these ideas by defining an extended proposal distribution that incorporates randomness in MC score estimates. In particular, we adopt an IS or AIS-based MC approach that also provides an unbiased estimate of the normalization constant, a property that, in principle, can be achieved by the rejection sampling approach of he2024zeroth as well. This property will be leveraged in the next part to construct intermediate targets for RDSMC.
3.2 Intermediate targets
Samples from the reverse diffusion proposal deviate from the desired target $\pi$ due to two sources of error: time discretization of the reverse SDE, and bias in score estimation. To correct these errors, we design a series of intermediate target distributions. We first characterize the optimal but intractable targets, and then develop practical approximations using byproducts of score estimation. Finally, we extend the intermediate targets to include auxiliary variables. Our construction aligns with the optimal targets at intermediate steps and recovers the desired marginal $\pi(x_{0})$ at the final step $t=0$.
Optimal intermediate targets.
The optimal intermediate target at step $t$ is the marginal distribution of the extended target $\pi(x_{0:T})$ in Eq.˜4,
$\pi(x_{t:T})=\int\pi(x_{0:T})\mathrm{d}x_{0:t-1}=\pi(x_{t})\prod_{i=t+1}^{T}\pi(x_{i}\mid x_{i-1}),$ (9)
as it leads to exact samples from $\pi(x_{0:T})$ when combined with locally optimal proposals [naesseth2019elements].
However, the marginal $\pi(x_{t})=\int\pi(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})\mathrm{d}x_{0}$ in Eq.˜9 involves an intractable integral, except in the final step $t=0$, where $\pi(x_{0})$ is known up to a normalization constant.
Marginal estimation.
For $t>0$, we approximate the marginal $\pi(x_{t})$ by $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ using byproducts of MC score estimates, where we recall $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} is the estimation randomness. The key insight is that $Z\pi(x_{t})$ is the normalization constant of the posterior $\pi(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$, that is,
$\int\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})\mathrm{d}x_{0}$ $=Z\int\pi(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})\mathrm{d}x_{0}=Z\pi(x_{t}).$
Hence, we can re-use the posterior inference procedure in MC score estimation to obtain an unbiased estimate of the desired marginal $\pi(x_{t})$ (up to a factor of $Z$).
For example, IS or AIS-based score estimation (Algorithms˜2 and 3) generates importance weights $\{w^{(m)}\}_{m=1}^{M}$ targeting $\pi(x_{0}\mid x_{t})$ for a fixed $x_{t}$; we then obtain an unbiased marginal estimate
$\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ $\coloneqq\frac{1}{M}\sum_{m=1}^{M}w^{(m)},\textrm{ with }\mathbb{E}_{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})}\left[\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\right]=Z\pi(x_{t}).$ (10)
At the final step $t=0$, we set $\hat{\pi}(x_{0},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0}})\coloneqq\tilde{\pi}(x_{0})=Z\pi(x_{0})$ as no approximation is needed.
Extended intermediate targets.
Finally, we incorporate the auxiliary variables $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} by setting their targets to match their sampling distributions and define the extended intermediate target as
$\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})$ $\coloneqq\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})\prod_{i=t}^{T-1}\pi(x_{i+1}\mid x_{i})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{i+1}}\mid x_{i+1})$ (11)
for $t=0,\cdots,T-1$, and $\gamma_{T}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})\coloneqq\hat{\pi}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})$.
The structure in Eq.˜11 mirrors that of the optimal targets in Eq.˜9, replacing the intractable marginal $\pi(x_{t})$ with an unbiased estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ (up to a factor of $Z$) and accounting for auxiliary randomness $\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}$}. Moreover, we make the following two observations.
Observation 1.
The final marginal target at $t=0$ matches the desired $\pi(x_{0})$ with no approximations,
$\begin{split}\gamma_{0}(x_{0})&=\int\gamma_{0}(x_{0:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}})\mathrm{d}x_{1:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}}\\ &=\int\hat{\pi}(x_{0},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{0}}\mid x_{0})\prod_{i=1}^{T}\pi(x_{i}\mid x_{i-1})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{i}}\mid x_{i})\mathrm{d}x_{1:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}}\\ &=\tilde{\pi}(x_{0})\propto\pi(x_{0}),\end{split}$ (12)
as $\hat{\pi}(x_{0},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0}})=\tilde{\pi}(x_{0})$ by construction. This implies our final SMC target is correctly specified.
Observation 2.
Marginalizing out $\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}$} in $\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})$ recovers the optimal target $\pi(x_{t:T}),\forall t$,
$\begin{split}\int\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})\mathrm{d}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}&=\int\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})\prod_{i=t}^{T-1}\pi(x_{i+1}\mid x_{i})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{i+1}}\mid x_{i+1})\mathrm{d}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}\\ &=\mathbb{E}_{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})}\left[\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\right]\,\prod_{i=t}^{T-1}\pi(x_{i+1}\mid x_{i})\\ &=Z\pi(x_{t})\,\prod_{i=t}^{T-1}\pi(x_{i+1}\mid x_{i})\propto\pi(x_{t:T}),\end{split}$ (13)
where the last equality follows from the unbiasedness of marginal estimates, as in Eq.˜10. This result shows that our intermediate targets $\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})$ are aligned with the optimal $\pi(x_{t:T})$, despite the approximation in $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, a property known as the exact approximation [andrieu2010particle, naesseth2019elements].
While the generic SMC framework remains valid under alternative choices of intermediate targets, provided that $\gamma_{0}(x_{0})\propto\pi(x_{0})$, our construction offers a balance between theoretical correctness and practical efficiency. The marginal estimates $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ in our intermediate targets from Eq.˜11 act as twisting or look-ahead functions [naesseth2019elements, Chapter 3] that incorporate future information $\pi(x_{0})$ into intermediate steps. Together, Observations˜1 and 2 ensure that our intermediate targets provide effective guidance throughout the sampling process while recovering the desired target in the end.
3.3 RDSMC: Algorithm and theoretical guarantee
We now derive weighting functions, the final component of RDSMC, which guides the resampling of particles. We then present the complete algorithm and establish its theoretical guarantees.
Input: Unnormalized target $\tilde{\pi}(x_{0})$, number of particles $N$, discretization steps $T$ (with step size $\delta=1/T$), diffusion schedule $\{\alpha_{t},\sigma_{t},f_{t},g_{t}\}$, base distribution $q(x_{T})$, and additional inputs for score and marginal estimation $\eta$
Output: Weighted samples $\{x_{0}^{(i)},w_{0}^{(i)}\}_{i=1}^{N}$ and normalization constant estimate $\hat{Z}$
1
2for $i\leftarrow 1$ to NN do
3   Sample $x_{T}^{(i)}\sim q(x_{T})$;
[…]
5   Compute weight: $w_{T}^{(i)}\leftarrow\frac{\hat{\pi}_{T}^{(i)}}{q(x_{T}^{(i)})}$;
6
7
8for $t\leftarrow T-1$ to 0 do
9   Resample $\{x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{(i)},s_{t+1}^{(i)},\hat{\pi}_{t+1}^{(i)}\}_{i=1}^{N}$ according to weights $\{w_{t+1}^{(i)}\}_{i=1}^{N}$ ;
10 for $i\leftarrow 1$ to NN do
11     Sample $x_{t}^{(i)}\sim q(x_{t}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{(i)}):=\mathcal{N}\left(x_{t}\mid x_{t+1}^{(i)}-\left[f_{t+1}x_{t+1}^{(i)}-g_{t+1}^{2}s_{t+1}^{(i)}\right]\delta,g_{t+1}^{2}\delta\right)$;
12     if t > 0 then
13        Compute score and marginal estimates $s_{t}^{(i)},\hat{\pi}_{t}^{(i)}\leftarrow$ Algorithm 3 $(\pi,\alpha_{t},\sigma_{t},x_{t}^{(i)},\eta)$;
14    else
15        Compute exact marginal $\hat{\pi}_{0}^{(i)}\leftarrow\tilde{\pi}(x_{0}^{(i)})$;
16
17    Compute weight: $w_{t}^{(i)}\leftarrow\frac{\hat{\pi}_{t}^{(i)}\,\pi(x_{t+1}^{(i)}\mid x_{t}^{(i)})}{\hat{\pi}_{t+1}^{(i)}\,q(x_{t}^{(i)}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{(i)})}$;
18
19
Compute normalization constant estimate $\hat{Z}\leftarrow\prod_{t=0}^{T}\frac{1}{N}\sum_{i=1}^{N}w_{t}^{(i)}$.
Algorithm 1 Reverse Diffusion Sequential Monte Carlo (RDSMC)
Weighting functions.
Following the standard SMC framework, we define the weighting functions on the extended space $\{x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}\}$ for $t=T-1,\cdots,0$ as ratios of intermediate targets and proposals
$w_{t}$ $\coloneqq\frac{\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})}{\gamma_{t+1}(x_{t+1:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1:T}})\,q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})}.$ (14)
Substituting the expressions for $\gamma_{t}$ from Eq.˜11 and canceling out common terms yields
$w_{t}$ $=\frac{\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\pi(x_{t+1}|x_{t})}{\hat{\pi}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}.$ (15)
At the initial step $T$, the weight is $w_{T}\coloneqq\frac{\gamma_{T}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})}{q(x_{T})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})}=\frac{\hat{\pi}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})}{q(x_{T})}$ recalling that $\gamma_{T}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})=\hat{\pi}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})$ by construction.
Notably, although the intermediate targets and proposals involve auxiliary sampling distributions $q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})$ for score estimation, the weighting functions $w_{t}$ in Eq.˜15 and $w_{T}$ do not depend on their evaluation. Hence, a generic MC sampler can be used while still retaining computable weights, as long as it produces tractable estimates of the score and marginal densities.
The RDSMC algorithm and theoretical guarantees.
We summarize the RDSMC algorithm in Algorithm˜1. It initializes $N$ particles from a tractable base distribution. Subsequently, the particles are propagated through a reverse diffusion-based proposal, followed by weighting and resampling using intermediate targets. These steps are enabled by running an inner-level MC estimation of the score function and the marginal density. In this work, we use AIS (Algorithm˜3) while other methods can be incorporated as well (see §˜B.1). The final output of RDSMC is a weighted set of samples approximating the target $\pi(x_{0})$, along with an estimate of the normalization constant $Z$.
RDSMC can be viewed as an adaptation of the nested SMC framework [naesseth2015nested, naesseth2019high]. The inner-level MC estimation involves a sampling procedure (e.g. AIS) targeting the intractable posterior $\pi(x_{0}\mid x_{t})$, which is then used to assign proper weights [naesseth2019elements, Chapter 4.3] to proposed samples $x_{t}$. Consequently, RDSMC inherits the unbiasedness and asymptotic exactness guarantees of nested SMC [naesseth2015nested, naesseth2019elements].
Theorem 1 (Informal).
Under regularity conditions, the RDSMC algorithm provides a consistent estimator of the target distribution $\pi(x_{0})$ as particle size $N\to\infty$ and an unbiased estimator of the normalization constant $Z$ for any $N\geq 1$.
We provide the formal statement and proof in Appendix˜C.
Theorem˜1 shows that RDSMC effectively wraps a diffusion MC-style proposal within an SMC framework, mitigating its bias as the number of particles increases. Moreover, it provides an unbiased estimate of the normalization constant, a property absent in existing diffusion MC methods [huang2023reverse, grenioux2024stochastic, he2024zeroth].
The bias correction mechanism of RDSMC arises from two key aspects: (1) the final target of RDSMC is explicitly constructed to match the desired target $\pi(x_{0})$, regardless of the discretization scheme or marginal density approximation (Observation˜1); and (2) errors in score estimation affect only the proposal steps, which are accounted for in the weighting functions. Therefore, the weighting and resampling steps in the outer SMC loop asymptotically correct for proposal bias, ensuring convergence to the final target $\pi(x_{0})$ in the limit of many particles.
4 Related works
Diffusion MC samplers.
huang2023reverse develop the RDMC sampler based on the reverse diffusion process using Langevin-style MCMC to estimate the score function. They initialize the sampler at some $\tau_{T}<1$ via a nested Langevin procedure. grenioux2024stochastic propose a similar SLIPS sampler with a signal-to-noise-ratio (SNR)-adjusted discretization scheme, and provide further guidance on choosing the initial sampling time $\tau_{T}$ under suitable conditions. In contrast, we use AIS-based score estimators to enable marginal estimation and start sampling from the base distribution at $\tau_{T}=1$.
Alternatively, he2024zeroth explore rejection sampling for score estimation, improving the sampling efficiency in low-dimensional settings, which may be incorporated into our framework as well.
While these methods provide certain theoretical guarantees, we offer an orthogonal means to improve their accuracy by increasing the number of particles. RDSMC can be viewed as a generic SMC wrapper around these diffusion-MC samplers (given suitable score estimators), enabling consistent sampling and unbiased estimation of the normalization constant.
Training-based diffusion samplers
In addition to developing MC estimates of the score function, another line of work trains a neural network directly using unnormalized target information. Variational approaches achieve this goal by minimizing the divergence between forward and reverse processes [zhang2021path, berner2022optimal, vargas2023denoising, vargas2023transport, richter2023improved], while akhound2024iterated, ouyang2024bnem exploit various identities of the score function or a related energy function. Similar to their training-free MC-based counterparts, these methods are prone to discretization and score approximation errors.
Closest to our work is phillips2024particle that also use SMC to correct proposal bias. Their method requires training neural networks, while ours is training-free. Moreover, their neural network-based intermediate targets incur approximation errors except at the last step, while ours reflect the true target marginals at each SMC iteration. See an empirical comparison in §˜E.5.3.
he2025no show that many training-based methods incur heavy computational overhead compared to classical sampling methods and rely on special network preconditioning. For this reason, we limit our empirical comparison to training-free samplers. Nonetheless, an interesting direction for future work is to combine our method with training-based approaches to further enhance performance.
SMC for conditional generation from diffusion models.
Beyond classical sampling tasks, SMC has been applied to conditional generation for pre-trained diffusion models [trippe2022diffusion, wu2023practical, cardoso2023monte, singhal2025general]. While these works also combine SMC and diffusion dynamics, a key distinction is the target distribution: they sample from an existing diffusion model prior tilted with a reward function, whereas our setting involves sampling from any (unnormalized) target distribution. As a result, the way diffusion models are used is different, and we require distinct designs for proposals and intermediate targets in SMC.
Nevertheless, RDSMC uses the estimated marginal densities in Eq.˜11 as twisting functions to improve SMC sampling efficiency. This strategy is conceptually is similar to that of wu2023practical, though in a different setting; see Appendix˜D for further discussion.
5 Experiments
We evaluate RDSMC on a range of synthetic and real-world target distributions, comparing it to SMC [del2006sequential], AIS [neal2001annealed], SMS [saremi2023chain], RDMC [huang2023reverse], and SLIPS [grenioux2024stochastic]. AIS and SMC operate over a series of geometric interpolations between the target and a Gaussian proposal with MCMC transitions. SMS samples from the target by sequentially denoising equally noised measurements using a Langevin procedure combined with estimated scores. RDMC and SLIPS are described in §˜4.
RDSMC uses a variance-preserving diffusion schedule. To evaluate the effectiveness of our intermediate targets and the importance of resampling, we include two ablations: (i) RDSMC (Proposal), which samples directly from the reverse diffusion proposal in Eq.˜8 without any weighting or resampling, and (ii) RDSMC (IS), which applies a final IS correction to samples from RDSMC (Proposal).
Baseline methods follow the implementation of grenioux2024stochastic. Notably, the information about the target variance (or an estimation) is provided to guide the initialization of the baselines SMC, AIS, SMS and SLIPS. In contrast, our method does not make use of this extra information.
Unless otherwise specified, we use $T=100$ discretization steps for RDSMC and its variants, and $T=1,024$ steps for other methods. We generate $N=4,096$ final samples for all methods, and tune their hyperparameters assuming access to either a validation dataset or an oracle metric.
We provide ablation studies controlling for the discretization steps, the total running time and comparable hyperparameter settings in §˜E.5, where we observe RDSMC still remains competitive and, in some cases, superior. All experiment details are included in Appendix˜E.
5.1 Bi-modal gaussian mixtures
(a) $d=2$: histogram along 1st dimension. While RDSMC’s proposal covers the modes, the SMC procedure is crucial for producing calibrated weights.
(b) Estimation bias of weight ratio $w_{1}/(w_{1}+w_{2})$ and log-normalization constant $\log Z$ versus dimension $d$. Results are averaged over 5 seeds with error bars showing one standard error. Note that SMS, RDMC, and SLIPS do not provide estimate of $\log Z$. For both metrics, the bias increases with $d$ for all methods, while RDSMC outperforms the baselines for most cases.
Figure 1: Bi-modal Gaussian mixture study.
We first study bi-modal Gaussian mixtures with an imbalanced weight ratio of $w_{1}/(w_{1}+w_{2})=0.1$ for varying dimensions $d$. The estimated ratio is obtained by assigning samples to the mode with the highest posterior probability. For each method we select the hyperparameters based on the lowest estimation bias of the weight ratio.
In Figure˜1(a), we compare the marginal histogram of samples from RDSMC, RDSMC (Proposal), and the true target in $d=2$. While samples from RDSMC (Proposal) cover both modes, their relative weights are overly balanced. In contrast, RDSMC recovers the calibrated mode weights, indicating the importance of the error correction mechanism by SMC.
Figure˜1(b) shows the estimation bias of the weight ratio (left), and that of the log-normalization constant $\log Z$ (right). Note that only RDSMC, RDSMC(IS), AIS and SMC provide estimates for $\log Z$. We observe that RDSMC consistently outperforms other methods in both metrics across dimensions. Moreover, RDSMC (Proposal) exhibits consistently high weight ratio estimation bias. While RDSMC (IS) reduces this bias, it still underperforms the full RDSMC procedure, highlighting the effectiveness of intermediate resampling guided by our intermediate target distributions.
In high dimensions, all methods exhibit some degree of mode collapse. RDSMC primarily samples from the dominant mode, resulting in average weight ratio estimation biases of 0.09 and 0.08 for $d=32$ and $64$, respectively. In contrast, AIS and SMC completely collapse to a single mode, which varies across runs, yielding degenerate estimated ratios of 0 or 1 and an average bias of around 0.4.
5.2 Rings and Funnel distributions
Algorithm Rings ($d=2$) Funnel ($d=10$)
Radius TVD $(\downarrow)$ $\log Z$ Bias $(\downarrow)$ Sliced KSD $(\downarrow)$ $\log Z$ Bias $(\downarrow)$
AIS 0.10 $\pm$ 0.00 0.05 $\pm$ 0.00 0.07 $\pm$ 0.00 0.28 $\pm$ 0.01
SMC 0.10 $\pm$ 0.00 0.05 $\pm$ 0.00 0.07 $\pm$ 0.00 0.28 $\pm$ 0.01
SMS 0.24 $\pm$ 0.01 N/A 0.15 $\pm$ 0.00 N/A
RDMC 0.37 $\pm$ 0.00 N/A 0.13 $\pm$ 0.00 N/A
SLIPS 0.19 $\pm$ 0.00 N/A 0.06 $\pm$ 0.00 N/A
RDSMC 0.13 $\pm$ 0.01 0.03 $\pm$ 0.00 0.11 $\pm$ 0.00 0.28 $\pm$ 0.10
RDSMC(IS) 0.15 $\pm$ 0.01 0.02 $\pm$ 0.01 0.33 $\pm$ 0.03 1.61 $\pm$ 0.14
RDSMC(Proposal) 0.09 $\pm$ 0.00 N/A 0.32 $\pm$ 0.03 N/A
Table 1: Results on Rings and Funnel (mean $\pm$ standard error over 5 seeds). Bold indicates 95% confidence interval overlap with the best average result. RDSMC has the lowest $\log Z$ estimation bias for both targets. On Rings, RDSMC(Proposal) has the lowest radius TVD, followed by AIS, SMC, and RDSMC. On Funnel, SLIPS has the lowest sliced KSD, followed by AIS, SMC, and RDSMC.
We present results on two additional synthetic targets. Rings, introduced by grenioux2024stochastic, is a 2-dimensional distribution constructed via an inverse polar parameterization: the angular component is uniformly distributed, while the radius component follows a 4-mode Gaussian mixture. Funnel [neal2003slice], is a 10-dimensional “funnel"-shaped distribution.
For Rings, we assess sample quality using total variational distance in the radius component (Radius TVD). For Funnel, use sliced Kolmogorov-Smirnov distance (Sliced KSD) to evaluate sample quality following grenioux2024stochastic. As both targets admit tractable normalization constants, we also report the estimation bias for methods that compute them. For each method, we select the hyperparameters with lowest Raidus TVD on a heldout validation set for Rings and lowest Sliced KSD for Funnel.
As shown in Table˜1, RDSMC achieves lower or comparable estimation bias of the log-normalization constant relative to AIS and SMC on both targets. On the lower-dimensional Rings target, RDSMC (Proposal) has the lowest radius TVD, where AIS, SMC and RDSMC closely match. On the more challenging Funnel target, SLIPS achieves the lowest sliced KSD; while RDSMC performs slightly worse, it greatly outperforms its IS and Proposal variants.
5.3 Bayesian logistic regression
Test LL $(\uparrow)$ Credit ($d=25$) Cancer ($d=31$) Ionosphere ($d=35$) Sonar ($d=61$)
AIS $\mathbf{-122.73\pm 0.51}$ $60.45\pm 0.31$ $\mathbf{-86.37\pm 0.10}$ $110.11\pm 0.06$
SMC $123.17\pm 0.05$ $60.28\pm 0.11$ $\mathbf{-86.37\pm 0.10}$ $110.11\pm 0.06$
SMS $527.79\pm 0.85$ $215.64\pm 0.66$ $202.56\pm 0.16$ $275.44\pm 0.31$
RDMC $388.24\pm 1.75$ $182.81\pm 0.43$ $108.67\pm 0.09$ $128.29\pm 0.03$
SLIPS $\mathbf{-121.79\pm 0.04}$ $\mathbf{-56.26\pm 0.08}$ $\mathbf{-85.07\pm 0.07}$ $\mathbf{-102.39\pm 0.03}$
RDSMC $\mathbf{-124.00\pm 1.96}$ $62.23\pm 1.96$ $\mathbf{-87.72\pm 1.75}$ $\mathbf{-101.52\pm 1.84}$
RDSMC(IS) $144.38\pm 4.63$ $82.47\pm 7.78$ $\mathbf{-84.90\pm 2.84}$ $\mathbf{-110.57\pm 4.54}$
RDSMC(Proposal) $606.03\pm 1.26$ $246.86\pm 0.97$ $92.62\pm 0.04$ $134.72\pm 0.13$
Table 2: Bayesian logistic regression with test log-likelihood (mean $\pm$ standard error) averaged over 5 seeds. Bold indicates 95% confidence interval overlap with that of the best average result. SLIPS achieves the best overall performance, while RDSMC matches or closely approaches it. In contrast, RDSMC(Proposal) performs noticeably worse, highlighting the effectiveness of the SMC correction.
Finally, we evaluate inference performance on Bayesian logistic regression models using four datasets. Credit and Cancer [nishihara2014parallel] involve predicting credit risk and breast cancer recurrence, while Ionosphere [sigillito1989classification] and Sonar [gorman1988analysis] focus on classifying radar and sonar signals, respectively. The inference is performed on $60\%$ of each dataset, leaving 20% for validation and 20% for testing. Each method is tuned using the validation log-likelihood estimate. We report the test log-likelihood in Table˜2.
RDSMC achieves the best or near-best performance across datasets, closely matching SLIPS, which performs the best overall. This comparison suggests that SLIPS’s time discretization and initialization strategy may offer complementary benefits despite lacking systematic error correction. RDSMC (IS) underperforms on Credit and Cancer and exhibits higher variance than RDSMC, while RDSMC (Proposal) performs the worst, again highlighting the importance of the SMC mechanism. Compared to baseline methods, RDSMC shows greater variance, likely due to auxiliary randomness in its nested structure, suggesting a direction for future improvement.
6 Discussion
This paper presents a training-free and diffusion-based SMC sampler, Reverse Diffusion Sequential Monte Carlo (RDSMC), for sampling from unnormalized distributions. By leveraging reverse diffusion dynamics as proposals, we devise informative intermediate targets to correct systematic errors within a principled SMC framework. These components rely on Monte Carlo score estimation, without requiring additional training. Our algorithm provides asymptotically exact samples from the target distribution and a finite-sample unbiased estimate of the normalization constant. Empirical results on both synthetic and Bayesian inference tasks demonstrate competitive or superior performance compared to existing diffusion-based and classical geometric annealing-based samplers.
Limitations and future work.
As a nested SMC procedure, RDSMC introduces auxiliary variance. Such variance may be mitigated by enhancing autocorrelation within the MC estimation step [e.g. deligiannidis2018correlated]. In high-dimensional Gaussian mixture experiments, we observe a degree of mode collapse, an issue also seen in other samplers. Potential remedies include partial resampling [martino2016weighting], or using more informed proposals. Our method also relies on oracle metrics for hyperparameter tuning, which aligns with existing works; however, developing a more automated strategy remains an important future direction. Finally, to further improve performance, we can incorporate SNR-adapted discretization schemes and informative initializations, as explored by grenioux2024stochastic, as well as combine our approach with training-based diffusion samplers.
Appendix A Background: diffusion models and diffusion-based samplers
A.1 Diffusion models
We review common diffusion schedules used in the literature.
Variance-preserving schedule.
In Variance-Preserving (VP) schedule [ho2020denoising], the coefficients of the forward SDE in Eq.˜1 are given by
$f(t)=-\frac{1}{2}b(t),\quad g(t)=\sqrt{b(t)},$ (16)
where $b(t)$ is a time-dependent function
$b(t)=b_{\min}+t(b_{\max}-b_{\min}),$
for some constants $b_{\min}\approx 0\ll b_{\max}$. For a large value of $b_{\max}$, the reference distribution is approximately standard normal, $p_{1}(x)\approx\mathcal{N}(x\mid 0,1)$.
By Ito’s calculus (see e.g. sarkka2019applied), one can show that the forward transition density $\pi(x_{s}\mid x_{t})$ for $1\geq s>t\geq 0$ is given by
$\pi(x_{s}\mid x_{t})=\mathcal{N}\left(x_{s}\mid\frac{\alpha(s)}{\alpha(t)}x_{t},\left(1-\frac{\alpha^{2}(s)}{\alpha^{2}(t)}\right)\mathbb{I}\right),$
where
$\alpha(t)=\exp\left(-\frac{1}{2}\int_{0}^{t}b(s)ds\right),\quad\sigma(t)^{2}=1-\exp\left(-\int_{0}^{t}b(s)ds\right).$
Variance-exploding schedule.
In Variance-Exploding (VE) schedule [song2020score], the forward SDE coefficients in Eq.˜1 are given by
$f(t)=0,\qquad g(t)=\sqrt{\frac{\mathrm{d}\sigma^{2}(t)}{\mathrm{d}t}},$
where $\sigma(t)$ is a non-negative and monotone increasing function with $\sigma(0)=0$. Commonly used forms of $\sigma(t)$ include polynomial functions, such as linear, quadratic or exponential schedules. When $\sigma(1)$ is large, the reference distribution can be effectively approximated as $\pi_{1}(x_{1})\approx\mathcal{N}(x_{1}\mid 0,\sigma^{2}(1))$.
Similarly to the VP case, one can derive the following parameters of the forward transition kernel with $\alpha(t)=1$ and $\sigma(t)$ is the given non-negative and monotone function.
A.2 Score identities used for diffusion MC samplers
In addition to the DSI from Eq.˜3, one may consider the following identities to form alternative MC estimates of the score, given MC samples of $\pi_{t}(x_{0}\mid x_{t})$.
Target score identity (TSI)
TSI utilizes known structural properties of the target density, such as estimated gradients, facilitating potentially more accurate and lower-variance score estimations at low-noise settings. Formally, the score function of the noisy target distribution $\pi_{t}(x)$ under the TSI framework satisfies
$\nabla_{x}\log\pi_{t}(x)$ $=\frac{1}{\alpha(t)}\int\nabla_{x}\log\pi_{t}(x_{0})\pi_{t}(x_{0}\mid x)dx_{0}.$ (17)
TSI and its related alternative have been used in several recent studies to improve score estimation accuracy [debortoli2024targetscorematching, phillips2024particle, akhound2024iterated, he2025featfreeenergyestimators].
Mixed score identity (MSI)
To balance the advantages of DSI and TSI across different noise regimes, MSI combines both identities through a convex interpolation. Specifically, MSI favors TSI in the low-noise regime and DSI in the high-noise regime, leveraging both strengths. Formally, MSI is a convex combination of TSI and DSI with coefficients $\frac{\alpha(t)^{2}}{\alpha(t)^{2}+\sigma(t)^{2}}$ and $\frac{\sigma(t)^{2}}{\alpha(t)^{2}+\sigma(t)^{2}}$:
$\nabla_{x}\log\pi_{t}(x)$ $=\frac{1}{\alpha(t)^{2}+\sigma(t)^{2}}\int\Big(\alpha(t)\big(x_{0}+\nabla_{x}\log\pi_{t}(x_{0})\big)-x(t)\Big)\pi_{t}(x_{0}\mid x)dx_{0}.$ (18)
he2024training leverage MSI to train an implicit generative model that produces samples in a single step, demonstrating improved sample quality and efficiency.
Appendix B Method details
In this section, we discuss several details of RDSMC, including score and marginal estimation strategies in §˜B.1, the computational complexity in §˜B.2, and implementation guidelines in §˜B.3.
B.1 Score and marginal estimation
In this subsection, we present several score and marginal estimators that can be used within RDSMC, which are based on IS, AIS, SMC and rejection sampling. In our main experiments we use AIS-based estimators.
B.1.1 IS-based estimator
[…]
Output: Score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ and marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$.
1 for $m\leftarrow 1$ to $M$ do
2   Sample ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\sim q(\cdot\mid x_{t})$
3   Compute importance weights $w^{(m)}\leftarrow\frac{\tilde{\pi}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})\mathcal{N}(x_{t}\mid\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}},\sigma_{t}^{2}I)}{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})}$
4
5Denote auxiliary variable ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\leftarrow\{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\}_{m=1}^{M}$
6 Compute score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\sum_{m=1}^{M}\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{M}w^{(m^{\prime})}}\cdot\frac{\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}-x_{t}}{\sigma_{t}^{2}}$
Compute marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{M}\sum_{m=1}^{M}w^{(m)}$.
Algorithm 2 Importance Sampling (IS)-based score and marginal estimator
The IS-based estimator is summarized in Algorithm˜2, which complements the illustration in §˜3.1.
huang2023reverse consider a similar importance weighted score estimator (see their Eq 5 in Section 3.3) where the proposal $q(\cdot\mid x_{t})=\mathcal{N}(\cdot\mid x_{t}/\alpha_{t},\sigma_{t}^{2}/\alpha_{t}^{2}\mathbb{I})$ is obtained by reversing the forward noising likelihood. While the IS-based score estimator can be inefficient when $t$ is large, huang2023reverse leverages Unadjusted Langevin Algorithm (ULA) with IS estimator for initialization; And when $t$ is close to 0, they are able to quickly obtain rough score estimates via the IS approach.
B.1.2 AIS and SMC-based estimator
[…]
Output: Score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ and marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$
1
2Set annealing schedule $\beta_{k}\leftarrow k/n$ for $k=0,\dots,n$
3 Set initial target $\nu_{1}(x_{0})\leftarrow q(x_{0}\mid x_{t})^{1-\beta_{1}}\tilde{\pi}(x_{0})^{\beta_{1}}\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})^{\beta_{1}}$
4 for $m\leftarrow 1$ to $M$ do
5   Draw initial sample ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(1,m)}}\sim q(\cdot\mid x_{t})$
6   Initialize weight $w^{(m)}\leftarrow\frac{\nu_{1}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(1,m)}})}{\nu_{0}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(1,m)}})}$
7
8for $k\leftarrow 2$ to $n$ do
9   Set the current target $\nu_{k}(x_{0})\propto q(x_{0}\mid x_{t})^{1-\beta_{k}}\tilde{\pi}(x_{0})^{\beta_{k}}\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})^{\beta_{k}}$
10 for $m\leftarrow 1$ to $M$ do
11      Update sample $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k,m)}$} using MCMC transition targeting $\nu_{k}$ starting from $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k-1,m)}$}
12      Update weight $w^{(m)}\leftarrow w^{(m)}\frac{\nu_{k}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k,m)}})}{\nu_{k-1}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k,m)}})}$
13
14
15Denote auxiliary variable ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\leftarrow\{\{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k,m)}}\}_{m=1}^{M}\}_{k=1}^{n}$
16 Compute normalized weights $\tilde{w}^{(m)}\leftarrow\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{M}w^{(m^{\prime})}}$
17 Compute score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\sum_{m=1}^{M}\tilde{w}^{(m)}\frac{\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(k,m)}}-x_{t}}{\sigma_{t}^{2}}$
Compute marginal density estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{M}\sum_{m=1}^{M}w^{(m)}$
Algorithm 3 Annealed Importance Sampling (AIS)-based Score and Marginal Estimator
Algorithm˜3 presents an AIS-based approach [neal2001annealed] to construct the score and marginal estimates. AIS introduces a sequence of intermediate unnormalized distributions, smoothly bridging a tractable initial proposal distribution $q(x_{0}\mid x_{t})$ to the denoising posterior $\pi(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$. Under a geometric annealing schedule, intermediate targets are $\nu_{k}(x_{0})\propto q(x_{0}\mid x_{t})^{1-\beta_{k}}\pi(x_{0}\mid x_{t})^{\beta_{k}}$ where $\beta_{k}=k/n$ for $k=0,\cdots,n$.
Algorithm˜3 begins by drawing a set of particles from the proposal distribution $q(x_{0}\mid x_{t})$. Then it gradually transitions these particles by running MCMC with a kernel invariant to the intermediate targets $\nu_{k+1}(x_{0})$ for $k=0,\cdots,n-1$. In this work we mainly consider Metropolis-Hasting adjusted Langevin dynamics [MALA, roberts1996exponential] and Hamiltonian Monte Carlo [HMC, neal2012mcmc]. At each intermediate step, the algorithm updates the importance weights of each particle based on the ratio of successive intermediate targets.
The final weighted set of particles provides a consistent estimate of the desired posterior, which is utilized to construct a score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, where $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} denotes all the randomness in this procedure. Additionally, these weights provide an unbiased estimate of marginal density $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, i.e. $\mathbb{E}_{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})}\left[\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\right]=Z\pi(x_{t})$, where $q(x_{t}\mid x_{t})$ is the sampling distribution of $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} given $x_{t}$.
SMC adaption of Algorithm˜3.
We can adapt Algorithm˜3 to an SMC-based procedure by incorporating the resampling mechanism. In this case, the weight update in Algorithm˜3 is modified to the incremental weight computation:
$w^{(k,m)}\leftarrow\frac{\nu_{k+1}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})}{\nu_{k}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})}$
and the initial weight in Algorithm˜3 is denoted by $w^{(1,m)}$, for $m=1,\cdots,M$. With this adjustment, we resample particles $\{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\}_{m=1}^{M}$ in the beginning of each iteration $k$ based on the previous weights $\{w^{(k-1,m)}\}_{m=1}^{M}$ for $k=2,\cdots,n$. The marginal estimate in Algorithm˜3 becomes
$\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{n}\sum_{k=1}^{n}\frac{1}{M}\sum_{m=1}^{M}w^{(k,m)}.$
The other aspect of Algorithm˜3 remains the same.
B.1.3 Rejection sampling (RS)-based estimator
[…]
Output: Score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ and marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$.
1
[…]
3 while $m\leq M$ do
4   Sample proposal ${\color[rgb]{0.3125,0.46875,0.62890625}v_{i}}\sim q(\cdot\mid x_{t})$
5   Compute acceptance ratio $r\leftarrow\frac{\tilde{\pi}({\color[rgb]{0.3125,0.46875,0.62890625}v_{i}})\,\mathcal{N}(x_{t}\mid\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}v_{i}},\sigma_{t}^{2}I)}{C\,q({\color[rgb]{0.3125,0.46875,0.62890625}v_{i}})}$
6   Draw ${\color[rgb]{0.3125,0.46875,0.62890625}v_{i}}\sim\mathrm{Uniform}(0,1)$
7 if $v<r$ then
8      Accept sample ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\leftarrow{\color[rgb]{0.3125,0.46875,0.62890625}v_{i}}$
9    $m\leftarrow m+1$
10
11 $i\leftarrow i+1$
12
[…]
14 Denote auxiliary variable ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\leftarrow\{{\color[rgb]{0.3125,0.46875,0.62890625}v_{i}},{\color[rgb]{0.3125,0.46875,0.62890625}v_{i}}\}_{i=1}^{M_{\textrm{total}}}$
15 Compute score estimate: $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{M}\sum_{m=1}^{M}\frac{\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}-x_{t}}{\sigma_{t}^{2}}$
16 Compute marginal estimate: $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow C\frac{M}{M_{\textrm{total}}}$
Algorithm 4 Rejection Sampling (RS)–based score and marginal estimator
In Algorithm˜4, we present a rejection sampling (RS)-based approach, building on prior work of he2024zeroth which propose an RS-based score estimator and develop theoretical guarantees for the resulting diffusion MC sampler. We extend their approach to additionally provide an unbiased marginal density estimate alongside the original score estimate. In principle, this procedure can be used within RDSMC (Algorithm˜1) as a replacement for the AIS scheme (Algorithm˜3). Future work may explore this direction particularly in low-dimensional settings, as he2024zeroth demonstrate advantages over other score estimators in a non-SMC context.
Notably, rejection sampling requires access to a constant $C$ that upper bounds the ratio of the unnormalized target to the proposal. he2024zeroth use Newton’s method to find a proxy of $C$.
B.1.4 Proposal choices in IS/AIS/SMC/RS-based score estimators.
Algorithms˜2, 3 and 4 require specifying an (initial) proposal distribution $q(x_{0}\mid x_{t})$. The choice of this proposal is flexible and should ideally approximate the true posterior $\pi(x_{0}\mid x_{t})$; the closer it is, the more effective the resulting estimator.
In Algorithms˜2, 3 and 4, the default initial proposal is obtained by reversing the forward noising likelihood, suggested by prior work [huang2023reverse, akhound2024iterated, he2024training]. This proposal is effective when the likelihood signal is strong—i.e., when $t$ is close to 0—but becomes less efficient as $t$ increases and the signal weakens.
Alternative strategies include using a Gaussian proposal centered at the current $x_{t}$. If additional information about the target distribution is available—such as its mean or covariance—it can be incorporated into the proposal. For instance, one may use a Laplace approximation of the target to construct an approximate posterior, which can be used as a sensible proposal.
B.1.5 Score estimator based on other score identities
In Eq.˜5 of the main paper, we derive the score estimator based on the DSI in Eq.˜3. Alternative score estimators can be derived based on other score identities.
Based on TSI in Eq.˜17, we obtain another estimate:
$\nabla_{x_{t}}\log p(x_{t})$ $\approx s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}):=\sum_{m=1}^{M}\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{M}w^{(m^{\prime})}}\cdot\frac{\nabla_{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}}\log q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})}{\alpha_{t}},$ (19)
where $\{w^{(m)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{(m)}\}_{m=1}^{M}$ is a weighted system approximating the posterior $\pi(x_{0}\mid x_{t})$ using an IS/AIS/SMC procedure.
Or, given MSI in Eq.˜18, we have:
$\nabla_{x_{t}}\log p(x_{t})$ $\approx s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}):=\sum_{m=1}^{M}\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{M}w^{(m^{\prime})}}\cdot\Big(\alpha_{t}\big({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}+\nabla_{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}}\log q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})\big)-x_{t}\Big).$ (20)
We primarily consider DSI and TSI-based estimators in our experiments.
B.2 Computational complexity
The total computational cost of RDSMC scales linearly with the number of diffusion discretization steps $T$, the number of particles $N$, and the cost of each score estimation step.
Specifically, for AIS-based score estimator as described in Algorithm˜3, the score estimation cost scales linearly with the number of importance samples $M$, the number of AIS transition steps $n$, and the number of MCMC steps per transition $l$. The vanilla IS-based score estimator in Algorithm˜2 is a special case with $n=1$ and $l=1$. Therefore the cost of score estimation per SMC particle per time step $\mathcal{O}(Mnl)$. Consequently, the overall complexity for RDSMC is $\mathcal{O}(NT\cdot Mnl)$.
Among these factors, the operations involving discretization steps $T$, AIS transition steps $n$, and MCMC steps $l$ must be processed sequentially due to their time-dependent structure. In contrast, computations over the particles $N$ and the importance samples $M$ are embarrassingly parallel and can be efficiently distributed across parallel compute resources.
B.3 Implementation guidelines
We discuss practical implementation guidelines for RDSMC, including the resampling strategy in §˜B.3.1, score clipping in §˜B.3.2 and hyperparameter choices in §˜B.3.3.
B.3.1 SMC resampling strategy
While the resampling step in RDSMC is crucial for steering particles toward high target probability region, it introduces extra variance in the short term. We discuss several strategies to reduce this added variance.
First, we use systematic resampling, which introduces correlations among resampling indices [naesseth2019elements, Chapter 2.2.1].
We also consider start resampling at a later stage of the sampling trajectory. At early steps, the Gaussian likelihood provides weaker conditional information, making posterior inference more challenging [huang2023reverse, grenioux2024stochastic]. Consequently, both the score estimates and marginal density estimates can be significantly biased or exhibit high variance, especially under limited compute budgets. As a result, the intermediate target distributions in early steps may not be sufficiently informative to enable effective resampling. To mitigate this issue, we delay resampling until a designated step $t_{\mathrm{start\text{-}resampling}}<T$, which we treat as a tunable hyperparameter.
For timestep $t<t_{\mathrm{start\text{-}resampling}}$, we adopt an effective sample size (ESS)-based criterion to decide whether resampling should be performed at each step – a common strategy in SMC literature [naesseth2019elements, Chapter 2.2.2]. The ESS at step $t$ is computed as
$ESS_{t}$ $:=\frac{\left(\sum_{i=1}^{N}w_{t}^{(i)}\right)^{2}}{\sum_{i=1}^{N}\left(w_{t}^{(i)}\right)^{2}}.$ (21)
We trigger resampling only when the normalized ESS, computed as $ESS_{t}/N$,is below a threshold $\kappa_{\mathrm{\text{ess}}}$, another hyperparameter which we fix at 0.3 throughout experiments.
When resampling is skipped (i.e. for $t>t_{\mathrm{start\text{-}resampling}}$ or when the normalized ESS is above $\kappa_{\mathrm{\text{ess}}}$), we must adjust the computation of the intermediate weights to ensure the unbiasedness of the log normalization constant estimate [naesseth2019elements, Chapter 2.2.3]. For iterations where resampling is not performed, the weight update in Algorithm˜1 of Algorithm˜1 is replaced by
$w_{t}^{(i)}$ $\leftarrow\frac{\bar{w}_{t+1}^{(i)}}{1/N}\cdot\frac{\hat{\pi}_{t}^{(i)}p(x_{t+1}^{(i)}\mid x_{t}^{(i)})}{\hat{\pi}_{t+1}^{(i)}q(x_{t}^{(i)}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{(i)})}$ (22)
where $\{\bar{w}_{t+1}^{(i)}\}_{i=1}^{N}$ are the normalized weights in the previous iteration.
B.3.2 Score clipping
To improve numerical stability, one may consider clipping the score estimates within a pre-specified range. Note that score clipping does not affect our theoretical guarantees provided that regularity conditions hold, as the score estimate is only part of the proposal mechanism.
As observed in prior work [noble2024learned], score estimates produced by the TSI estimator defined in Eq.˜17 can exhibit high variance or have large magnitudes, especially during early steps. akhound2024iterated apply score norm clipping when training diffusion samplers where the loss function is based on the TSI. Inspired by their work, we also clip the norm of the score estimates to a maximum of 20 when we use the TSI-based score estimates. In contrast, we do not apply clipping when using the DSI-based score estimator.
B.3.3 Hyperparameter choices
The main hyperparameters of RDSMC are:
Diffusion-related hyperparameters.
  • Diffusion parameterization: we use the VP diffusion schedule described in Eq.˜16 in this work.
  • Time discretization strategy: we use a uniform time discretization grid with $T=100$ steps, though more informative strategies can be incorporated—such as the SNR-adapted scheme proposed by grenioux2024stochastic.
MC score estimation-related hyperparameters.
  • Score estimator type: we use AIS-based score estimator (Algorithm˜3), with vanilla IS (Algorithm˜2) as a special case. We mainly leverage the DSI (Eq.˜5) to estimate the score in most experiments, while we also use the TSI (Eq.˜19) in Bayesian logistic regression tasks.
  • MCMC transition kernel: MALA or HMC.
  • $M$: Number of importance samples used in Monte Carlo score estimation
  • $n$: Number of intermediate distributions, i.e. number of AIS steps; for IS $n=1$
  • $\delta_{\text{mcmc}}$: Step size used in the MCMC transition kernel (not used in IS). We manually tune this parameter in the main experiments while we use adaptive step size in ablation studies; see discussion below.
  • $l$ : Number of MCMC steps per AIS transition (not used in IS). We use $l=1$ in all experiments.
SMC-related hyperparameters.
  • $t_{\mathrm{start\text{-}resampling}}$: Starting time for resampling within the SMC trajectory. No resampling is performed for $t/T>t_{\mathrm{start\text{-}resampling}}$.
  • $\kappa_{\mathrm{\text{ess}}}$: Normalized effective sample size threshold triggering resampling. When $\kappa_{\mathrm{\text{ess}}}=0$, no resampling is applied throughout, which corresponds to RDSMC (IS) if a final-step weighting is applied and RDSMC (Proposal) otherwise. We set $\kappa_{\mathrm{\text{ess}}}=0.3$ in all experiments for RDSMC.
In the main experiments, we choose score estimator and MCMC kernel types by a coarse grid search, and tune $M,n,\delta_{\text{mcmc}}$ and $t_{\mathrm{start\text{-}resampling}}$ in a range of values, using validation metrics. The hyperparameter values are summarized Table˜3; see details in §˜E.1.2.
In the ablation experiments in §˜E.5, we fix most hyperparameters to reduce the degrees of freedom. We use $T=100,M=100,n=50,m=1,\kappa_{\mathrm{\text{ess}}}=0.3$, and AIS-based score estimator leveraging the DSI, LG as the MCMC kernel. We also use an adaptive MCMC step size $\delta_{\text{mcmc}}$, as in grenioux2024stochastic. When the acceptance rate is too high ($76\%$) we increase the step size by $3\%$, when it is $74\%$, we decrease step size by $3\%$, and otherwise we keep the current step size. Only an initial step size value needs to be specified and then it is adpatively adjusted during the sampling process. The only hyperparameter in RDSMC that requires tuning is $t_{\mathrm{start\text{-}resampling}}$ , which we vary in $\{0,0.1,\cdots,0.9,1.0\}$. We observe consistent performance and therefore in practice we would recommend users to start with something similar to this configuration.
Appendix C Theoretical guarantees of RDSMC
We formally state the theorem providing sufficient conditions for asymptotic accuracy of RDSMC and unbiasedness of the estimate of the normalization constant. Our proof strategy for asymptotic accuracy is adapted from wu2023practical and the proof for unbiasedness follows the techniques presented in naesseth2019elements.
Theorem 2.
Assume
  1. (a)
    the first marginal estimate $\hat{\pi}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}),$ and the ratios of subsequent ratio $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})/\hat{\pi}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$ are positive and bounded,
  2. (b)
    the score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ is bounded, and $x_{t}$ has compact support,
  3. (c)
    variance increases in the forward noising kernel, i.e. for each $t,\ g_{t+1}^{2}>g_{t}^{2}.$
Then the RDSMC algorithm described in Algorithm˜1 provides a consistent estimator of the target distribution $\pi(x_{0})$ as particle size $N\to\infty$ and an unbiased estimator of the normalization constant $Z$ for any $N\geq 1$. More specifically,
  1. (a)
    Let $\mathbb{P}_{N}=\sum_{i=1}^{N}w_{0}^{(i)}\delta_{x_{0}^{(i)}}$ for weighted particles $\{(x_{0}^{(i)},w_{0}^{(i)})\}_{i=1}^{N}$ returned by Algorithm˜1 with $N$ particles. $\mathbb{P}_{N}$ converges setwise to $\pi(x_{0})$ with probability one, that is for every set $A,$ $\lim_{N\rightarrow\infty}\mathbb{P}_{N}(A)=\int_{A}\pi(x_{0})dx_{0}.$
  2. (b)
    For the estimate for the normalization constant $\hat{Z}$ returned by Algorithm˜1, we have $\mathbb{E}\hat{Z}=Z$.
The assumptions of Theorem˜2 are typically satisfied in standard implementations of SMC samplers:
  • Assumption (a) can be satisfied by the clipping of the marginal estimates $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ within the range $[c,C]$ for some constants $c,C>0$.
  • Assumption (b) by clipping scores estimates as well. Additionally, the compactness of the state space $x_{t}$ is a standard and commonly adopted assumption in the SMC literature.
  • Assumption (c) is mild and typically satisfied in practice. In both VE and VP noising schedules commonly used in diffusion models, the variance of the forward kernel increases with time.
Proof of Theorem˜2:
We first prove the asymptotic exactness of our sampler and then prove the unbiasedness of the normalization constant estimate.
First, Theorem˜3 characterizes a set of conditions under which SMC algorithms converge. We restate this result below in our own notation.
Theorem 3 (chopin2020introduction – Proposition 11.4).
Let $\{(x_{0}^{(i)},w_{0}^{(i)})\}_{i=1}^{N}$ be the particles and weights returned at the last iteration of a sequential Monte Carlo algorithm with $N$ particles using multinomial resampling. If each weighting function $w_{t}$ is positive and bounded, then for every bounded, $\nu_{0}$- measurable function $\phi$ of $x_{t}$
$\lim_{N\rightarrow\infty}\sum_{i=1}^{N}w_{0}^{(i)}\phi(x_{0}^{(i)})=\int\phi(x_{0})\nu_{0}(x_{0})dx_{0}.$
with probability one.
An immediate consequence of Theorem˜3 is the setwise convergence of the discrete measures. To apply Theorem˜3to show asymptotic accuracy, it is sufficient to show that the weights at each step are upper bounded, as they are defined through probabilities and hence are positive. Since there is a finite number of steps $T$, it suffices to show that each $w_{t}$ is bounded.
The initial weight is defined by
$w_{T}=\frac{\hat{\pi}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})}{q(x_{T})},$
which is bounded by Assumption (a).
The intermediate weights are defined by Eq.˜14 such that
$w_{t}=\frac{\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\,\pi(x_{t+1}\mid x_{t})}{\hat{\pi}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\,q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}.$
We first decompose the weights by
$\log w_{t}=\log\frac{\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})}{\hat{\pi}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}+\log\frac{\pi(x_{t+1}\mid x_{t})}{q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}.$
The boundedness of $\log\frac{\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})}{\hat{\pi}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}$ is guaranteed by Assumption (a). We now show the boundedness of $\log\frac{\pi(x_{t+1}\mid x_{t})}{q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}$. First, note that
$\pi(x_{t+1}\mid x_{t})=\mathcal{N}(x_{t+1}\mid x_{t}+f_{t}x_{t}\delta,g_{t}^{2}\delta),$
and
$q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})=\mathcal{N}(x_{t}\mid x_{t+1}-\left[f_{t+1}x_{t+1}-g_{t+1}^{2}s(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\right]\delta,g_{t+1}^{2}\delta).$
Let $\hat{\mu}_{p}=x_{t+1}-f_{t}x_{t}\delta\ \text{ and }\ \hat{\mu}_{q}=x_{t+1}-\left[f_{t+1}x_{t+1}-g_{t+1}^{2}s(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\right]\delta.$ We can show that $\|\hat{\mu}_{p}-\hat{\mu}_{q}\|=\|f_{t+1}x_{t+1}-f_{t}x_{t}-g_{t+1}^{2}s(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\|\delta$ is bounded , which follows from Assumption (b). The log-ratio then simplifies as
$\log\frac{\pi(x_{t+1}\mid x_{t})}{q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}$ (23)
$=\log\frac{|2\pi g_{t}^{2}I|^{-1/2}\exp\{-(2g_{t}^{2})^{-1}\|x_{t}-\hat{\mu}_{p}\|^{2}\}}{|2\pi g_{t+1}^{2}I|^{-1/2}\exp\{-(2g_{t+1}^{2})^{-1}\|x_{t}-\hat{\mu}_{q}\|^{2}\}}$ (24)
$\textrm{Rearrange \ and\ let \ }C=\log|2\pi g_{t}^{2}I|^{-1/2}/|2\pi g_{t+1}^{2}I|^{-1/2}$ (25)
$=-\frac{1}{2}\left[g_{t}^{-2}\|x_{t}-\hat{\mu}_{p}\|^{2}-g_{t+1}^{-2}\|x_{t}-\hat{\mu}_{q}\|^{2}\right]+C$ (26)
$\textrm{Expand and rearrange \ }\|x_{t}-\hat{\mu}_{q}\|^{2}=\|x_{t}-\hat{\mu}_{p}\|^{2}+2\langle\hat{\mu}_{p}-\hat{\mu}_{q},x_{t}-\hat{\mu}_{p}\rangle+\|\hat{\mu}_{p}-\hat{\mu}_{q}\|^{2}$ (27)
$=-\frac{1}{2}\Big[(g_{t}^{-2}-g_{t+1}^{-2})\|x_{t}-\hat{\mu}_{p}\|^{2}-2g_{t+1}^{-2}\langle\hat{\mu}_{p}-\hat{\mu}_{q},x_{t}-\hat{\mu}_{p}\rangle-g_{t+1}^{-2}\|\hat{\mu}_{p}-\hat{\mu}_{q}\|^{2}\Big]+C$ (28)
$\textrm{Let \ }C^{\prime}{=}C+\frac{1}{2}g_{t+1}^{2}\|\hat{\mu}_{p}-\hat{\mu}_{q}\|^{2}\textrm{\ and rearrange. Note that $\|\hat{\mu}_{p}-\hat{\mu}_{q}\|^{2}{<}\infty$}$ (29)
$=-\frac{1}{2}(g_{t}^{-2}-g_{t+1}^{-2})\|x_{t}-\hat{\mu}_{p}\|^{2}+g_{t+1}^{-2}\langle\hat{\mu}_{p}-\hat{\mu}_{q},x_{t}-\hat{\mu}_{p}\rangle+C^{\prime}$ (30)
By Cauchy-Schwarz, (31)
$\leq-\frac{1}{2}(g_{t}^{-2}-g_{t+1}^{-2})\|x_{t}-\hat{\mu}_{p}\|^{2}+g_{t+1}^{-2}\|\hat{\mu}_{p}-\hat{\mu}_{q}\|\cdot\|x_{t}-\hat{\mu}_{p}\|+C^{\prime}$ (32)
[…]
$\leq\frac{1}{2}\frac{(g_{t+1}^{-2}\|\hat{\mu}_{p}-\hat{\mu}_{q}\|)^{2}}{g_{t}^{-2}-g_{t+1}^{-2}}+C^{\prime}$ (34)
$=\frac{g_{t+1}^{-4}}{2(g_{t}^{-2}-g_{t+1}^{-2})}\|\hat{\mu}_{p}-\hat{\mu}_{q}\|^{2}+C^{\prime}$ (35)
$\leq C^{\prime\prime}.$ (36)
The final line follows from the boundedness of $\|\hat{\mu}_{p}-\hat{\mu}_{q}\|$. The above derivation shows that each $w_{t}$ is bounded, which concludes the proof for the asymptotic exactness of our sampler.
Then we show the unbiasedness of the normalization constant estimate $\hat{Z}$. Let $a^{i}_{t}$ be the index of $x_{t}^{(i)}$ and $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(i)}$} before resampling. With that, $\hat{Z}$ is defined by
$\hat{Z}=\prod_{t=0}^{T}\frac{1}{N}\sum_{i=1}^{N}w_{t}^{(i)},\quad\text{where }w_{t}^{(i)}=\frac{\hat{\pi}_{t}^{(i)}\pi(x_{t+1}^{(i)}\mid x_{t}^{a^{i}_{t}})}{\hat{\pi}_{t+1}^{(i)}q(x_{t}^{a^{i}_{t}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})},\ w_{T}^{(i)}=\frac{\hat{\pi}_{T}^{(i)}}{\hat{\pi}(x_{T}^{a^{i}_{t}})}.$
Denote the normalized weight by $\widetilde{w}_{t}^{(i)}=\frac{w_{t}^{(i)}}{\sum_{i=1}^{N}w_{t}^{(i)}}$. The distribution of all random variables generated by the SMC method is then given by
$Q(x_{0:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}},a_{0:T})$
$=\pi(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})p(x_{0:{T-1}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:{T-1}}},a_{0:T})$
$=\pi(x_{t}^{a_{T}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}})\pi(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T}^{a_{T}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}})\prod_{t=0}^{T-1}p(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t}^{a_{t}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}})q(x_{t}^{a_{t}}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}}\mid x_{t}^{a_{t}}),$
$=\pi(x_{t}^{a_{T}})\Big(\prod_{i=1}^{N}\frac{w_{T}^{(i)}}{\sum_{i=1}^{N}w_{T}^{(i)}}\Big)\prod_{i=1}^{N}\Big[\Big(\prod_{t=0}^{T-1}\frac{w_{t}^{(i)}}{\sum_{i=1}^{N}w_{t}^{(i)}}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})\Big)\Big]$
$\Big[q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}}\mid x_{T}^{a_{T}})\prod_{t=0}^{T-1}q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}}\mid x_{t}^{a_{t}})\Big].$
By $w_{t}^{(i)}=\frac{\hat{\pi}_{t}^{(i)}\pi(x_{t+1}^{(i)}|x_{t}^{a_{t}^{i}})}{\hat{\pi}_{t+1}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})}$ and $w_{T}^{(i)}=\frac{\hat{\pi}_{T}^{(i)}}{q(x_{T}^{a_{T}^{i}})}$, we have:
$Q(x_{0:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}},a_{0:T})$
$=\Big(\prod_{t=0}^{T}\sum_{i=1}^{N}w_{t}^{(i)}\Big)^{-1}\hat{\pi}_{0}^{(1)}\Big(\prod_{t=0}^{T-1}\pi(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})\Big)$
$\prod_{i=2}^{N}\hat{\pi}_{T}^{(i)}\prod_{i=2}^{N}\Big[\Big(\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})\Big)\Big]\Big[q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}}\mid x_{T}^{a_{T}})\prod_{t=0}^{T-1}q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}}\mid x_{t}^{a_{t}})\Big],$
where we plug in $w_{t}^{(1)}$ and rearrange terms.
By multiplying $\hat{Z}$, we have:
$\hat{Z}Q(x_{0:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}},a_{0:T})$ $=\hat{Z}\Big(\prod_{t=0}^{T}\sum_{i=1}^{N}w_{t}^{(i)}\Big)^{-1}\hat{\pi}_{0}^{(1)}\Big(\prod_{t=0}^{T-1}\pi(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})\Big)$
$\prod_{i=2}^{N}\Big[\hat{\pi}_{T}^{(i)}\Big(\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})\Big)\Big]\Big[q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}}\mid x_{T}^{a_{T}})\prod_{t=0}^{T-1}q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}}\mid x_{t}^{a_{t}})\Big]$
$=N^{-T}\hat{\pi}_{0}^{(1)}q(u_{0}^{a_{0}^{1}}\mid x_{0}^{a_{0}^{1}})\Big(\prod_{t=0}^{T-1}\pi(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{1}}\mid x_{t+1}^{a_{t+1}^{1}})\Big)$
$\prod_{i=2}^{N}\Big[\hat{\pi}_{T}^{(i)}q(u_{t}^{a^{i}_{T}}\mid x_{t}^{a^{i}_{T}})\Big(\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{(i)}}\mid x_{t+1}^{a_{t+1}^{(i)}})\Big)\Big],$
where the first equality follows from plugging in $Q$ and the second follows from plugging in $\hat{Z}$.
The expectation of the normalization constant estimate can be written as
[…]
$\prod_{i=2}^{N}\Big[\hat{\pi}_{T}^{(i)}q(u_{t}^{a^{i}_{T}}\mid x_{t}^{a^{i}_{T}})\Big(\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{(i)}}\mid x_{t+1}^{a_{t+1}^{(i)}})\Big)\Big]dx_{0:T}d{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}}$
$=$ $N^{-T}\sum_{{a_{0:T}^{1}}}\int\int\hat{\pi}_{0}^{(1)}q(u_{0}^{a_{0}^{1}}\mid x_{0}^{a_{0}^{1}})\Big(\prod_{t=0}^{T-1}\pi(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{1}}\mid x_{t+1}^{a_{t+1}^{1}})\Big)dx_{0:T}^{1}du_{0:T}^{1}$
$=$ $Z,$
where in the second equality, we marginalize the variables not involved in the particle $x_{0:T}^{1}$ and $x_{0:T}^{a_{0:T}^{1}}$. The final equality follows because we are averaging over $N^{T}$ possible cases of $a_{0:T}^{1}$ and all are equal to Z, which concludes the proof.
Appendix D Comparison to Twisted Diffusion Sampler
In this section, we discuss similarities and differences between RDSMC and Twisted Diffusion Sampler [TDS, wu2023practical], which is an SMC method for conditional generation from diffusion models.
Structurally, our method shares similarities with TDS, as both methods use some form of "look-ahead" or "twisting" functions [naesseth2019elements, Chapter 3] to improve the efficiency of SMC. These twisting functions introduce future information to intermediate states, promote promising particles through the weighting procedure, and improve the SMC sampling efficiency.
Moreover, the incremental weight expression of both TDS and RDSMC takes the same form (their Eq. (11) v.s. our Eq.˜15):
$w_{t}$ $:=\frac{\psi(x_{t})\gamma^{0}_{t}(x_{t:T})}{\psi(x_{t+1})\gamma^{0}_{t}(x_{t+1:T})q(x_{t}\mid x_{t+1})},$
where $\psi$ is the twisting function, $\gamma^{0}_{t}$ is the untwisted intermediate target and $q$ is the proposal. The product $\gamma_{t}(x_{t:T})\coloneqq\psi(x_{t})\gamma_{t}^{0}(x_{t:T})$ defines a series of twisted intermediate targets.
While the expressions of these components are method-specific, the structure of the weighting functions is shared between TDS our RDSMC and follows from the generic twisted SMC framework [naesseth2019elements, Chapter 3].
Importantly, this weight construction introduces twisting functions in a telescoping manner such that their approximation errors cancel out over the full trajectory. This property ensures that the final target is correct, and the resulting SMC procedure is asymptotically exact for both TDS and RDSMC.
However, we want to highlight several important differences between TDS and RDSMC:
  1. 1.
    Problem setting: The problem we address is fundamentally different from that of TDS. TDS is designed to sample from the conditional distribution $p_{\theta}(x_{0}\mid y)\propto p_{\theta}(x_{0})p(y\mid x_{0})$ given a pretrained diffusion model $p_{\theta}(x_{0})$ and a likelihood model $p(y\mid x_{0})$ with the observation $y$. As a comparison, our goal is to sample from an arbitrary distribution $\pi(x_{0})$, assuming access to its unnormalized density $\tilde{\pi}(x_{0})$.
  2. 2.
    Design of key components: Due to the difference in settings, our intermediate target distributions, proposal distributions, and twisting functions require different designs. In particular,
    • TDS’s twisting function is $p(y\mid\hat{x}_{\theta}(x_{t})$ (their Eq. (8)), the likelihood evaluated at the denoising prediction $\hat{x}_{\theta}(x_{t})$ by the pretrained model, which approximates the intractable likelihood $p_{\theta}(y\mid x_{t})$. This twisting function incorporates the observation $y$ into the weighting function (their Eq. (11)). Notably, when $t=0$, TDS’s twisting function recovers the exact likelihood $p(y\mid x_{0})$ and no approximation is involved.
      Given the twisting function, TDS’s intermediate target is $p\left(y\mid\hat{x}_{\theta}(x_{t})\right)p_{\theta}(x_{t:T})$ where $p_{\theta}(x_{t:T})$ is the distribution under the pretrained diffusion model. The proposal is obtained by adding a term that is the gradient of the twisting function, $\nabla_{x_{t}}\log p(y\mid\hat{x}_{\theta}(x_{t}))$, to the pretrained diffusion model’s transition kernel (see their Eq (9) and (11)).
    • In our case, RDSMC uses the marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ (Eq.˜10) as the twisting function, which approximates the exact marginal $\pi(x_{t})$. The marginal integrates out the future information in $\pi(x_{0})$, thereby encoding the target information into the intermediate weight.
      We define intermediate targets as $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\pi(x_{t:T})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}\mid x_{t:T})$ (a restatement of Eq.˜11); the product term $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\pi(x_{t:T})$ is structurally similar to TDS’s intermediate target, while we additionally account for MC estimation randomness $q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}\mid x_{t:T})$. We also utilize twisting function to inform the proposal, that is, we use the score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, which is connected to the twisting function / marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, to design the proposal in Eq.˜8.
  3. 3.
    Use of auxiliary variables: Our method introduces auxiliary randomness through the Monte Carlo estimation of the twisting function. Incorporating this randomness places RDSMC in the category of nested SMC methods [naesseth2015nested]. In contrast, TDS does not use auxiliary variables and operates more like a standard SMC algorithm, without requiring nested sampling.
Appendix E Experiment details
This section provides experimental details and additional results.
§˜E.1 describes the implementation and hyperparameter settings for RDSMC and all baseline methods. §˜E.2, E.3 and E.4 present supplementary information and additional results for the target distributions used in the main experiments. §˜E.5 reports additional experiments that control for hyperparameters, runtime, and other relevant factors to compare across methods, as well as an empirical comparison to phillips2024particle.
Compute resources.
All experiments were conducted on a single NVIDIA A100 GPU with 40 GB GPU memory, and accessed between 16 GB and 32 GB of CPU memory, depending on the task. RDSMC and baseline methods are all implemented in PyTorch [paszke2019pytorch].
E.1 Method and hyperparameter details
We first give an overview of baseline methods we benchmark RDSMC against in §˜E.1.1. We then describe key hyperparameter settings of our method and competing baselines for the main experiments in §˜E.1.2.
E.1.1 Summary of baseline methods
AIS [neal2001annealed].
The annealed importance sampling (AIS) algorithm defines a sequence of geometrically annealed distributions $\rho_{t}(x)$ for $t\in\{0,1,\cdots,T\}$ from an initial proposal $\rho_{0}(x)$ to the desired target $\rho_{T}(x)\propto\pi(x)$ as $\rho_{t}(x)\propto\rho_{0}(x)^{1-\beta_{t}}\pi(x)^{\beta_{t}}$ where $\{\beta_{t}\}_{t=0}^{T}$ is an increasing linear schedule with each $\beta_{t}=t/T$. AIS begins by sampling from the initial proposal $\rho_{0}(x)$ and proceeds through a sequence of MCMC transitions, each targeting the intermediate distribution $\rho_{t}(x)$. The resulting weighted samples provide a consistent approximation to the target distribution $\pi(x)$. The complexity of AIS is $\mathcal{O}(NTn)$ for $N$ final samples, $T$ annealing steps, and $n$ MCMC transitions per step.
SMC [del2006sequential].
Sequential Monte Carlo (SMC) operates on the same sequence of annealed distributions as AIS but differs by performing resampling at certain iterations. In our implementation, we consider both constant resampling at each step and an adaptive strategy based on an ESS threshold $\kappa_{\mathrm{\text{ess}}}$. The complexity of SMC is the same as that of AIS.
RDMC [huang2023reverse].
Reverse diffusion Monte Carlo (RDMC) builds on the time‐reversed Ornstein–Uhlenbeck diffusion:
$dY_{t}=\bigl\{Y_{t}+2\,\nabla\log p_{t_{\mathrm{\text{init}}}-t}(Y_{t})\bigr\}\,dt+\sqrt{2}\,dB_{t},\quad Y_{0}\sim\pi_{t_{\mathrm{\text{init}}}}$
where $p_{s}(y)=\int\mathcal{N}(y;\,\mathrm{e}^{-s}x,\,(1-\mathrm{e}^{-2s})I)\,\pi(dx)$. By Tweedie’s formula, the intractable score $\nabla\log p_{t_{\mathrm{\text{init}}}-t}(Y_{t})$ can be written in terms of the posterior denoiser:
$u_{t}(y)\;=\;\int x\,q_{t}(x\mid y)\,dx,\qquad q_{t}(x\mid y)\propto\pi(x)\,\mathcal{N}\bigl(x;\,\mathrm{e}^{t_{\mathrm{\text{init}}}-t}y,\,(\mathrm{e}^{2(t_{\mathrm{\text{init}}}-t)}-1)\,I\bigr),$
so that the SDE becomes
$dY_{t}=\big\{\tfrac{\mathrm{e}^{-2\,(t_{\mathrm{\text{init}}}-t)}+1}{\mathrm{e}^{-2\,(t_{\mathrm{\text{init}}}-t)}-1}\,Y_{t}+\tfrac{2\,\mathrm{e}^{-(t_{\mathrm{\text{init}}}-t)}}{1-\mathrm{e}^{-2\,(t_{\mathrm{\text{init}}}-t)}}\,u_{t}(Y_{t})\big\}dt+\sqrt{2}\,dB_{t}.$
Following the implementation of grenioux2024stochastic, RDMC discretizes the interval $0,t_{\mathrm{\text{init}}}]$ into $T$ steps. At initialization, RDMC employs a Langevin-within-Langevin scheme to sample from $\pi_{t_{\mathrm{\text{init}}}}$: using ULA where the score function is estimated based on DSI. In subsequent steps, RDMC simulates from the reverse SDE where the denoiser $u_{t}(y)$ is obtained by running MALA targeting $q_{t}(x\mid y)$.
The overall complexity of RDMC is $\mathcal{O}(NM(n_{\textrm{init}}+Tn))$ where $N$ is the number of final samples, $T$ is the number of iterative sampling steps, and the remaining factors are associated with score estimation: $M$ is the number of MCMC chains, while $n_{\textrm{init}}$ and $n$ denote the number of MCMC transition steps in the initial and subsequent stages, respectively.
SLIPS [grenioux2024stochastic].
Stochastic Localization via Iterative Posterior Sampling (SLIPS) relies on a stochastic observation process defined as:
$Y_{t}=\alpha(t)X+\sigma W_{t},$
where $(W_{t})_{t\geq 0}$ is a standard Brownian motion independent of $X\sim\pi$, and $\alpha(t)$ is a flexible denoising schedule function. To bypass the direct sampling requirement from $\pi$, SLIPS introduces a conditional denoiser defined by:
$u_{t}(y)=\int xq_{t}(x\mid y)dx,\quad q_{t}(x\mid y)\propto\pi(x)\mathcal{N}\left(\frac{y}{\alpha(t)},\frac{\sigma^{2}}{g(t)^{2}}I\right).$
The corresponding SDE governing this observation process is:
$dY_{t}=\dot{\alpha}(t)u_{t}(Y_{t})dt+\sigma dB_{t},$ (37)
where $(B_{t})_{t\geq 0}$ is another standard Brownian motion.
Similar to RDMC, the SLIPS algorithm approximates the dynamics defined by this SDE using an MCMC approach to estimate the conditional denoiser $u_{t}$ or a related score function:
  • The SLIPS initialization also involves a Langevin-within-Langevin procedure. Different from RDMC, SLIPS chooses the initialization time $t_{\mathrm{\text{init}}}$ such that both the marginal $\pi_{t_{\mathrm{\text{init}}}}(y)$ and the posterior $q_{t_{\mathrm{\text{init}}}}(x\mid y)$ are (approximately) log-concave , which they refer to as the “duality of log-concavity" assumption.
  • In the subsequent steps, SLIPS integrates the observation process where the denoiser $u_{t}(Y_{t})$ is estimated by MALA sampling from $q_{t}(x\mid y)$.
Similar to RDMC, SLIPS’ complexity is $\mathcal{O}(NM(n_{\textrm{init}}+Tn))$ where $N$ is the number of final samples, $T$ is the number of iterative sampling steps, and the remaining factors are associated with score estimation: $M$ is the number of MCMC chains, while $n_{\textrm{init}}$ and $n$ denote the number of MCMC transition steps in the initial and subsequent stages, respectively.
SMS [saremi2023chain].
For a target $\pi$, Sequential Multimeasurement Sampler (SMS) draws $M$ independent noisy observations at some noise scale $\sigma>0$:
$Y^{m}=X+\sigma Z^{m},\quad Z^{m}\sim\mathcal{N}(0,I),\quad m=1,\dots,M,$
with $X\sim\pi$. For any $m\in\{1,...,M\}$, the posterior density of $X$ given $y_{1:m}$ is
$q_{m}(x\mid y_{1:m})\;\propto\;\pi(x)\,\mathcal{N}\!\bigl(\,\bar{y}_{1:m},\,\tfrac{\sigma^{2}}{m}I\bigr),$
where $\bar{y}_{1:m}=(1/m)\sum_{i=1}^{m}y_{i}$. Its Bayes estimator
$u_{m}(y_{1:m})\;=\;\mathbb{E}[X\mid Y^{1:m}=y_{1:m}]$
serves as a non‐Markovian denoiser, and one shows that the law of $u_{m}(Y^{1:m})$ converges to $\pi$ in $W_{2}$ at rate $O(\sigma\sqrt{d/m})$ [saremi2023chain].
To sample from $\pi$, one first simulates the entire $M$-tuple $Y^{1:M}$ and then returns $u_{M}(Y^{1:M})$. They employ a ‘Once-At-A-Time’ strategy: draw $Y^{1}$ via ULA, then sequentially sample each $Y^{m}\mid Y^{1:m-1}$ using MCMC targeting the conditional density of $Y^{m}$. Under mild assumptions this sequence of targets becomes increasingly log-concave in $m$, but the initial draw of $Y^{1}$ can be as challenging as sampling from $\pi$ itself when $\sigma$ is large. Although one can estimate scores via IS or inner-loop posterior MCMC, numerical results show a steep degradation in performance unless $\sigma$ is carefully tuned—a manifestation of the same “duality of log-concavity” that forces a trade-off between ease of initialization and overall convergence as reported by grenioux2023onsampling.
The time complexity of SMS is $O(NTMn)$ for $N$ final samples, $T$ iterative sampling steps, $n$ MCMC transition steps per sampling step, and $M$ importance samples used for score estimation.
E.1.2 Implementation details and hyperparameter settings for the main experiments
Target $M$ $n$ $\delta_{\text{mcmc}}$ $t_{\mathrm{start\text{-}resampling}}$ score est. mcmc kernel
GMM ($d=2,4$) {10, 100} {1, 10} {1.0} {0.5, 0.8} DSI HMC
GMM ($d=8\text{ - }64$) {10, 100} { 1, 10, 50} {1.0} {0.5, 0.8} DSI HMC
Rings {100} {1, 10} {0.01, 0.05} {0.5} DSI MALA
Funnel {100} {1, 10, 50} {0.01, 0.05} {0.5, 0.8} DSI MALA
Logistic Reg. { 10, 100} {10, 50} {0.01, 0.005} {0.5, 0.8} TSI MALA
Table 3: Hyperparameter grid used for tuning RDSMC in the main experiments. We fix the number of sampling steps at $T=100$, set $l=1$ MCMC step per AIS transition, and use an ESS threshold of $\kappa_{\mathrm{\text{ess}}}=0.3$. Score estimator and MCMC kernel types are selected via a coarse grid search using validation metrics. RDSMC (IS) and RDSMC (Proposal) use the same grid with $\kappa_{\mathrm{\text{ess}}}=0$, excluding $t_{\mathrm{start\text{-}resampling}}$ as it does apply in these settings.
We summarize the hyperparameter values used to tune RDSMC and its variants for the main experiments in Table˜3, where these hyperparameters are explained in §˜B.3.3. The choices of score estimator and MCMC kernel types are determined by a coarse grid search using validation metrics; We primarily tune $M,n,\delta_{\text{mcmc}}$ and $t_{\mathrm{start\text{-}resampling}}$, as described in Table˜3.
Target AIS SMC SMS
$\delta_{\text{mcmc-init}}$ $\delta_{\text{mcmc-init}}$ $\kappa_{\mathrm{\text{ess}}}$ $\delta_{\text{mcmc}}$ $\delta_{\gamma}$
GMM {0.1, 0.5, 1.0} {0.1, 0.5, 1.0} {0.3, 1.0} {0.03, 1.0} {0.0625, 0.05}
Rings {0.01} {0.01} {0.3, 1.0} {0.03, 1.0} {0.0625, 0.05}
Funnel {0.01} {0.01} {0.3, 1.0} {0.03, 1.0} {0.0625, 0.05}
Logistic Reg. {0.01} {0.01} {0.3, 1.0} {0.03, 1.0} {0.0625, 0.05}
Target RDMC SLIPS
$\exp\{t_{\mathrm{\text{init}}}\}$ $t_{\mathrm{\text{final}}}$ $\epsilon$ $\delta_{\text{mcmc-init}}$
GMM {$0.95,0.9,0.8,0.7$ } {150, 300} {0.03, 0.05, 0.1, 0.2, 0.4} {0.1, 0.5, 1.0}
Rings {$0.95,0.9,0.8,0.7$ } {150, 300} {0.1, 0.2, 0.4, 1.0, 1.2} {1e-5}
Funnel {$0.95,0.9,0.8,0.7$ } {150, 300} {0.1, 0.2, 0.4, 1.0, 1.2} {1e-5}
Logistic Reg. {$0.95,0.9,0.8,0.7$ } {150} {0.03, 0.05, 0.1, 0.2, 0.4} {1e-5}
Table 4: Hyperparameter grid used for tuning baseline methods in the main experiments. AIS, SMC, RDMC, and SLIPS use $T=1,000$ sampling steps, and SMS uses $K=500$ noisy observations.
The hyperparameter search grid for baseline methods is provided in Table˜4. We largely follow the implementation of the official codebase of grenioux2024stochastic (available at https://github.com/h2o64/slips). For the bi-modal GMM experiments, we tune key hyperparameters within a similar range to those used for RDSMC. For the Rings, Funnel, and Bayesian logistic regression tasks, we adopt the settings reported by grenioux2024stochastic.
We use the following target variance information to initialize AIS, SMC, and SLIPS. Crucially, RDMC and our method do not make use of this information.
We restate Assumption A0 from grenioux2024stochastic.
Assumption 1.
(Log-concavity outside a compact). There exist $R>0$ and $\tau>0$ such that $\pi$ is the convolution of $\mu$ and $\mathcal{N}(0,\tau^{2}\,\mathbf{I}_{d})$, where $\mu$ is a distribution compactly supported on $\mathbb{B}=\mathbb{B}(\mathbf{m}_{\pi},R)$, i.e., $\mu(\mathbb{R}^{d}\setminus\mathbb{B})=0$.
The values of $R$ and $\tau$ that approximately or exactly satisfy this assumption for different targets are provided in the corresponding experiment subsections.
We next discuss the hyperparameter settings for each of the baseline methods.
AIS.
AIS uses $T=1,000$ annealing steps. Each transition step uses an MALA kernel with an adaptive step size, initialized at $\delta_{\text{mcmc-init}}$, to maintain the acceptance ratio at 75%. Additionally, MALA runs 4 parallel chains, each with 32 MCMC steps. The initial proposal $\rho_{0}(x)$ is set to $\mathcal{N}(x\mid 0,R^{2}+\tau^{2})$.
The only tunable parameter we consider is $\delta_{\text{mcmc-init}}$.
SMC.
SMC follows the same configuration as AIS, except it incorporates ESS-based resampling, introducing the normalized ESS threshold $\kappa_{\mathrm{\text{ess}}}$ as an additional hyperparameter (alongside $\delta_{\text{mcmc-init}}$).
RDMC.
RDMC uses $T=1,000$ discretization steps. To initialize $\pi_{t_{\mathrm{\text{init}}}}$ in RDMC, the Langevin-within-Langevin algorithm is simulated using 16 MCMC steps and 4 MCMC chains. The chains are initialized with an IS approximation of the posterior powered by 128 particles. The initial sample is drawn from $\mathcal{N}(0,(1-\exp\{-2t_{\mathrm{\text{init}}}\})\mathbb{I}_{d}\})$ and the initial step size is set to $(1-\exp\{-2t_{\mathrm{\text{init}}}\})/2$
Subsequent steps use a MALA kernel for posterior sampling with adaptive step size to maintain the acceptance ratio at 75%, 4 parallel chains and 32 MCMC steps. We drop the first 50 MCMC samples to ignore the warm-up period in the estimation.
The only tunable parameter we consider is $t_{\mathrm{\text{init}}}$.
SLIPS.
SLIPS follows a similar initialization and posterior sampling scheme as RDMC, also with $T=1,000$ discretization steps. However, instead of tuning the value of the initial sampling time $t_{\mathrm{\text{init}}}$, SLIPS makes use of the target variance information and sets $\pi_{t_{\mathrm{\text{init}}}}(x):=\mathcal{N}(x\mid 0,R^{2}/d+\tau^{2})$.
SLIPS also features a few algorithmic subtleties.
  • SLIPS uses a stochastic exponential integrator scheme to simulate from their observational SDE, where ours and RDMC consider the Euler Maruyama scheme.
  • SLIPS discretizes the SDE using evenly spaced points in the log-SNR space, while our method and RDMC perform discretization in the time domain.
  • SLIPS reuses the final MCMC samples from the previous step as the initialization for MCMC at the next step, promoting continuity across iterations.
  • At the final step, SLIPS outputs the estimated denoiser as an approximate sample from the target $\pi$ to align with standard stochastic localization literature. In contrast, our method and RDMC return a random sample generated from the approximated SDE dynamics.
We consider the following hyperparameters: (i) $t_{\mathrm{\text{final}}}$ the final integration time of Eq.˜37, (ii) $\epsilon$ the initial time to determine the log-SNR-based discretization schedule, and (iii) $\delta_{\text{mcmc-init}}$, the initial step size for the MALA kernel.
SMS.
We set the number of noisy observations $M$ in SMS to $\lfloor T/2\rfloor$ where $T=1,000$, and use as many MCMC steps per noise level as SLIPS or RDMC. This choice of $K$ ensures that the computational complexity of SMS is on par with other baseline methods. The Langevin steps are done using ULA as suggested by the authors.
We tune the step size $\delta_{\text{mcmc}}$ and the friction coefficient $\delta_{\gamma}$ using recommended values by the authors.
E.2 Bi-modal gaussian mixtures
Target.
We consider a Gaussian mixture model (GMM) with two components in a $d$-dimensional space. The component means $m_{1},m_{2}$ are randomly initialized from a uniform distribution over a box of width $80$, centered at the origin. Each component has diagonal covariance with all diagonal values equal to $\sigma^{2}=2\log 2$. The weights are fixed to $w_{1},w_{2}]=[0.1,0.9]$ regardless of dimension $d$. Formally, the target density is
$\pi(x)$ $=w_{1}\mathcal{N}(x\mid m_{1},\sigma^{2}\mathbb{I}_{d})+w_{2}\mathcal{N}(x\mid m_{2},\sigma^{2}\mathbb{I}_{d}).$ (38)
Metrics.
We compute the estimated weights by assigning samples to different components.
For a sample $x$, we assign it to the component $k\in\{1,2\}$ with highest posterior probability value, which is computed as
$p(x\textrm{ in component }k)=\frac{w_{k}\,\mathcal{N}(x\mid\mu_{k},\sigma^{2}\mathbb{I}_{d})}{\sum_{j=1}^{2}w_{j}\,\mathcal{N}(x\mid\mu_{j},\sigma^{2}\mathbb{I}_{d})}.$
The weight estimates are then formulated as $\hat{w}_{1}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{x_{i}\textrm{ in component }1\}$ and $\hat{w}_{2}=1-\hat{w}_{1}$.
The weight estimation ratio bias is computed as $|\hat{w}_{1}/(\hat{w}_{1}+\hat{w}_{2})-w_{1}/(w_{1}+w_{2})|=|\hat{w}_{1}-w_{1}|$.
The log normalization constant estimation bias is simply $|\hat{Z}-Z|$ where $\hat{Z}$ is the estimated value, for method that computes it.
Additional implementation details.
Assumption˜1 are satisfied for $R=\max(w_{1},w_{2})\|m_{1}-m_{2}\|=0.9\|m_{1}-m_{2}\|$, and $\tau=\sigma$. These values are used to initialize AIS, SMC, SMS, and SLIPS.
For RDSMC and its variants, we use DSI to estimate the score. The initial proposal distribution in Algorithm˜3 is set to $\mathcal{N}(x_{0}\mid x_{t}/\alpha_{t},\sigma_{t}^{2}/\alpha_{t}^{2})$.
Figure 2: Estimation bias of weight ratio $w_{1}/(w_{1}+w_{2})$ and log normalization constant $\log Z$, with hyperparameters selected to minimize the $\log Z$ estimation bias. We observe that all methods exhibit monotonic trends in both bias metrics.
Additional results.
The hyperparameters used for results in Figure˜1(b) are selected based on minimizing the estimation bias of weight ratio $w_{1}/(w_{1}+w_{2})$. Here, we also report results using hyperparameters tuned by minimizing the $\log Z$ bias in Figure˜2. Results for RDSMC (Proposal), SMS, RDMC, and SLIPS are not included as they do not provide normalization constant estimate. The trends are largely similar, except that the $\log Z$ estimation bias curve appears monotonic.
E.3 Rings and Funnel distributions
Targets.
The Rings distribution [grenioux2024stochastic] is defined via an inverse polar reparameterization of a base distribution $p_{z}$, which is factorized into two independent univariate marginals $p_{r}$ and $p_{\theta}$:
  • $p_{r}$ is a mixture of four Gaussian distributions, each with mean located at integer radial positions $i+1$ for $i\in\{0,1,2,3\}$ and variance $0.15^{2}$.
  • The angular component $p_{\theta}$ is uniformly distributed over $0,2\pi]$.
The Funnel distribution [neal2003slice] is characterized by a hierarchical density structure in a 10-dimensional space, where the first dimension $x_{1}$ is drawn from a Gaussian distribution with zero mean and variance $\sigma^{2}=9$. The subsequent dimensions $x_{2:10}$ are conditionally Gaussian, with variances exponentially dependent on the first dimension. We follow grenioux2024stochastic to use the following Funnel density
$\pi(x_{1:10})=\mathcal{N}(x_{1};0,\sigma^{2})\mathcal{N}(x_{2:10};0,\exp(x_{1})\mathbb{I}_{9}).$ (39)
Metrics.
The entropic regularized 2-Wasserstein distance between two distributions $\mu$ and $\nu$ is defined by
$W_{2,\varepsilon}(\mu,\nu)=\inf\{\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x_{1}-x_{0}\|^{2}\mathrm{d}\pi(x_{0},x_{1})-H(\pi):\pi_{0}=\mu,\pi_{1}=\nu\}^{1/2}\;,$ (40)
where $\varepsilon>0$ is a regularization hyper-parameter and $H(\pi)=-\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\log\pi(x_{0},x_{1})\mathrm{d}\pi(x_{0},x_{1})$ refers to as the entropy of $\pi$. We evaluate the quality of sampling with regularization $\varepsilon=0.05$ via POT library [flamary2021pot].
Total Variation Distance (TVD) measures the discrepancy between the true and model-generated distributions by comparing their marginal histograms within a bounded region.
Let $\hat{\mu}_{i}$ and $\hat{\nu}_{i}$ represent the probability mass in the $i$-th bin for the true and generated distributions, respectively. We compute the TVD as follows
$\text{TVD}=\frac{1}{2}\sum_{i}|\hat{\mu}_{i}-\hat{\nu}_{i}|,$
For the radius TVD, we generate the histogram of the radius samples using 256 uniform bins over the interval $[0,8]$. For the angle TVD, we generate the histogram of angle samples using 256 uniform bins over the interval $\pi,\pi]$.
The Kolmogorov-Smirnov distance (KSD) between two distributions $\mu$ and $\nu$ is defined by
$\operatorname{KSD}(\mu,\nu)=\sup_{x\in\mathbb{R}^{d}}\|F_{\mu}(x)-F_{\nu}(x)\|\;,$ (41)
where $F_{\mu}$ and $F_{\nu}$ denotes the cumulative distribution function of $\mu$ and $\nu$ respectively. We evaluate the 10-dimensional Funnel samples using the sliced version from [grenioux2023onsampling, Appendix D.1].
Additional implementation details.
For both targets, we use DSI to estimate the score. On Rings, we choose $R=4,\tau=0.15$ that satisfy Assumption˜1 to initialize AIS, SMC, SMS, and SLIPS; and on Funnel, we use $R=2.12,\tau=0.0$, following the implementation of grenioux2024stochastic.
On Rings, we set the initial proposal distribution in Algorithm˜3 for RDSMC and its variants to $\mathcal{N}(x_{0}\mid x_{t}/\alpha_{t},\sigma_{t}^{2}/\alpha_{t}^{2})$. On Funnel, the above choice would render numerical instability when $t$ is large. Hence we use $\mathcal{N}(x_{0}\mid x_{t},\sigma_{t}^{2}/\alpha_{t}^{2})$ as the proposal for the AIS procedure.
E.4 Bayesian logistic regression
Targets.
We consider four Bayesian logistic regression datasets: Credit and Cancer [nishihara2014parallel], Ionosphere [sigillito1989classification] and Sonar [gorman1988analysis].
  • Credit [nishihara2014parallel]: this dataset addresses binary classification of individuals as good or bad credit risks. It contains 1000 data points with 25 standardized features.
  • Cancer [nishihara2014parallel]: this dataset involves classifying recurrence events in breast cancer. It consists of 569 data points in a 31-dimensional standardized feature space.
  • Ionosphere [sigillito1989classification]: this dataset classifies radar signals passing through the ionosphere into good or bad categories. It features 351 data points with dimensionality $d=35$.
  • Sonar [gorman1988analysis]: this dataset differentiates sonar signals reflected from metal cylinders versus cylindrical rocks. It includes 208 data points with $d=61$ standardized features.
We consider a training dataset $\mathcal{D}=\{(x_{j},y_{j})\}_{j=1}^{M}$ where $x_{j}\in\mathbb{R}^{d}$ and $y_{j}\in\{0,1\}$ for all $j\in\{1,\ldots,M\}$. A Bayesian logistic regression model consists of the following
$p(w,b)$ $=\mathcal{N}(w;0,\mathbf{I}_{d})\mathcal{N}(b;0,(2.5)^{2})$ (42)
$p(y\mid x;w,b)$ $=\text{Bernoulli}(y;\sigma(x^{\top}w+b))$ (43)
where $w\in\mathbb{R}^{d}$ is a weight vector, $b\in\mathbb{R}$ is an intercept, and $\sigma$ is the sigmoid function.
The target distribution we are interested in is the posterior distribution
$p(w,b\mid\mathcal{D})\propto p(\mathcal{D}\mid w,b)p(w,b)=\prod_{j=1}^{M}p(y_{j}\mid x_{j};w,b)p(w,b).$ (44)
Metric.
Following grenioux2024stochastic, we use the following definition of the predictive log-likelihood given a test dataset $\mathcal{D}_{\text{test}}$:
$\frac{1}{N}\sum_{i=1}^{N}\sum_{(x,y)\in\mathcal{D}_{\text{test}}}\log p(w_{i},b_{i})+\log p(y\mid x;w_{i},b_{i})$ (45)
where $\{w_{i},b_{i}\}_{i=1}^{N}$ are posterior samples.
Additional implementation details.
On all tasks, we use $R=2.5,\tau=0.0$ in Assumption˜1 to initialize AIS, SMC, SMS, and SLIPS following the implementation of grenioux2024stochastic.
For RDSMC and its variants, we use TSI to estimate the score and clip the norm of the estimated score within 20 to improve numerical stability. We use $\mathcal{N}(x_{0}\mid x_{t}/\alpha_{t},\sigma_{t}^{2}/\alpha_{t}^{2})$ as the initial proposal distribution in the AIS-based score estimation procedure.
E.5 Ablation experiments
In our main experiments, baseline methods follow the implementation of grenioux2024stochastic which use more discretization steps $T$ compared to our method, and result in different runtimes. Additionally, AIS, SMC and SLIPS use target variance information while RDSMC does not.
In this section, we conduct ablation studies to control for these factors and evaluate performance under comparable computational budgets. We discuss the hyperparameter settings in §˜E.5.1. We present the results for the targets in the main experiments with the new hyperparameter settings in §˜E.5.2.
In addition, we include an empirical comparison to the Particle Denoising Diffusion Sampler [phillips2024particle] in §˜E.5.3.
E.5.1 Hyperparameter settings
To enable fair comparisons under similar computational budgets, we fix the compute budget for RDSMC and align the hyperparameters of SLIPS, RDMC, and SMS accordingly, since these methods have comparable computational complexity; however, algorithmic differences lead to variations in runtime. For AIS and SMC, we adjust their hyperparameters to achieve total runtimes comparable to that of RDSMC.
In addition, only SLIPS and SMS use target scalar variance to initialize their samplers, while other methods including RDSMC do not access this information.
We summarize hyperparameter settings for each method below.
RDSMC.
  • Use $T=100$ uniform discretization timesteps
  • Use AIS score estimator based on the DSI, and MALA as the MCMC kernel in AIS
  • Use an adaptive MCMC step size, similar to that used in the baseline SLIPS, AIS, and SMC methods (see description in §˜B.3.3). This adaptive strategy differs from the fixed step size employed in the main experiments which requires tuning. The initial MCMC step size is 1 for GMM targets and 0.05 for others.
  • Use $M=100$ AIS importance samples
  • Use $n=50$ AIS annealing steps
  • Use $m=1$ MCMC transition per AIS step
  • Use $\kappa_{\mathrm{\text{ess}}}=0.3$ for the ESS-based resampling threshold
  • The only tunable hyperparameter is the starting step for resampling, $t_{\mathrm{start\text{-}resampling}}$, which we vary in $\{0,0.1,0.2,\cdots,1.0\}$ unless otherwise specified. Recall that we apply resampling at step $t$ when $t\leq t_{\mathrm{start\text{-}resampling}}$ and the normalized ESS $\kappa_{\mathrm{\text{ess}}}$.
We treat RDSMC (IS) as a special case of RDSMC with $t_{\mathrm{start\text{-}resampling}}=0$, which is included within the generic RDSMC framework as part of hyperparameter tuning. The final hyperparameter values tuned by validation metrics are reported in Table˜5.
Target start resampling time ($t_{\mathrm{start\text{-}resampling}}$)
GMM $(d=2)$ 0.2
GMM $(d=4)$ 0.6
GMM $(d=8)$ 0.4
GMM $(d=16)$ 0.2
GMM $(d=32)$ 0.1
GMM $(d=64)$ 0.4
Rings 0.1
Funnel 1.0
Credit 0.9
Cancer 0.6
Ionosphere 0.7
Sonar 1.0
Table 5: Ablation experiments: final hyperparameter settings for RDSMC, selected by target-dependent validation metrics.
RDSMC (Proposal) uses the same configuration as RDSMC except that no resampling or weighting is applied throughout the sampling process.
Target # of annealing steps ($T$) # of MCMC transitions per step ($n$)
GMM ($d=2,4,8$) 100 100
GMM ($d=16$) 150 100
GMM ($d=32$) 240 100
GMM ($d=64$) 400 100
Rings 100 70
Funnel 100 100
Credit 1000 140
Cancer 1000 90
Ionosphere 1000 70
Sonar 1000 60
Table 6: Ablation experiments: final hyperparameter settings for AIS and SMC, tuned to achieve comparable runtime to RDSMC (with $T=100$ time steps, $M=100$ AIS importance samples, and $n=50$ AIS annealing steps) on a single NVIDIA A100 GPU (40 GB memory). While the $M$ importance samples in RDSMC can be processed in parallel, computation can still result in slower performance on a single device. We therefore increase $T$ and $n$ for AIS and SMC to match the runtime of RDSMC.
AIS and SMC.
To generate $N$ final samples, recall that the computational complexity of AIS and SMC is $\mathcal{O}(NTn)$ for $T$ annealing steps and $n$ MCMC transitions per step; in contrast, RDSMC’s complexity is $\mathcal{O}(NTMn)$ for $T$ discretization steps, $M$ AIS importance samples, $n$ AIS annealing steps (with a minor abuse of notation).
While the $M$ AIS importance samples in RDSMC can be processed in parallel, computation can still result in in slower runtime on a single device. Therefore, we adjust the number of annealing steps $T$ and the number of MCMC transitions per step $n$ to ensure that the total runtime on a single NVIDIA A100 GPU (40 GB memory) is comparable to that of RDSMC. The resulting hyperparameter values are provided in Table˜6.
Additionally, in contrast to the main experiments following grenioux2024stochastic, we do not use the target variance to initialize the initial proposal in AIS and SMC. Instead, we use the initial proposal $\mathcal{N}(0,1)$ which coincides with the base distribution of RDSMC. We use the same adaptive MCMC step size strategy as in RDSMC. For SMC, the ESS-based resampling threshold is set to 0.3.
Other aspects of AIS and SMC are the same as in the main experiments.
SLIPS and RDMC.
We match the computational budget of each score estimation step to that of RDSMC. Specifically, both SLIPS and RDMC use $T=100$ discretization steps; in every step, they use $M=100$ parallel MCMC chain, each with $n=50$ transitions, for score estimation. Consequently, the asymptotic complexity of SLIPS and RDMC (after the initialization stage) is the same as that of RDSMC, which is $\mathcal{O}(NTMn)$ to produce $N$ final samples.
However, because SLIPS and RDMC require additional computation for initialization and their score estimation relies on MCMC, whereas RDSMC starts from the base distribution and uses AIS for score estimation, the overall runtimes of these methods differ.
Other aspects of SLIPS and RDMC are the same as in the main experiments
SMS.
We use $T=100$ steps, $n=50$ MCMC transitions per sampling step and $M=100$ importance samples for score estimation. Other aspects of SMS are the same as in the main experiments.
E.5.2 Ablation experiment results
We revisit the targets studied in the main experiment using the new hyperparameter settings and summarize the results. Each experiment is run with 10 random seeds, and hyperparameters are selected using the corresponding validation metrics.
Runtime.
Figure 3: Ablation experiments: runtimes (in seconds) for all methods across different targets. Each blue bar represents the total wall-clock time averaged over 10 random seeds with error bar denoting 2 standard errors, measured on a single NVIDIA A100 GPU (40 GB). For RDMC and SLIPS, we also include the initialization time indicated by the orange bars. RDSMC(P) denotes RDSMC (Proposal). The GMM results are reported separately in the final panel of Figure˜4. AIS and SMC are configured to achieve runtimes comparable to RDSMC. While RDMC, SLIPS and SMS have the same asymptotic complexity of RDSMC, their actual runtimes differ due to algorithmic differences.
The overall runtimes for all methods across all targets are shown in Figure˜3, except for the GMM targets, which are presented in the final panel of Figure˜4.
We note that AIS and SMC are manually configured to achieve runtimes comparable to RDSMC. While RDMC, SLIPS and SMS have the same asymptotic complexity of RDSMC, their actual runtimes differ due to algorithmic differences. In GMM (with varying $d$), Rings, and Funnel, we observe shorter runtimes for RDMC, SLIPS, and SMS compared to RDSMC. In contrast, for Credit, Cancer, Sonar, and Ionosphere, SLIPS tends to have a longer runtime. Notably, the initialization phase of RDMC and SLIPS accounts for a substantial portion of their total runtime.
Bi-modal GMM.
Figure 4: Ablation experiments: estimation bias of weight ratio $w_{1}/(w_{1}+w_{2})$ (left) and log-normalization constant $\log Z$ (middle), and runtime (in seconds, right) versus dimension $d$. Results are averaged over 10 seeds with error bars showing one standard error. Note that SMS, RDMC, and SLIPS do not provide estimates of $\log Z$. For both metrics, the estimation bias increases with $d$ for all methods, while RDSMC consistently outperforms the baselines. The runtime for RDSMC (including the Proposal variant). SMC and AIS are configured for comparable runtime of RDSMC, whereas SLIPS, RDMC, and SMS exhibit shorter runtimes owing to algorithmic differences, despite having the same asymptotic complexity as RDSMC.
We report the results in Figure˜4. In the first two panels, we observe that for both the weight ratio $w_{1}/(w_{1}+w_{2})$ and the log normalization constant $\log Z$, RDSMC consistently outperforms the other methods. These results are consistent with the findings in the main experiments in Figure˜1, despite a different hyperparameter setting. The last panel summarizes the runtime comparison.
The optimal hyperparameters for each method are selected based on the lowest estimation bias of the weight ratio, consistent with the procedure used in the main experiments.
Rings and Funnel.
Algorithm Rings Funnel
Radius TVD $\downarrow$ log Z bias $\downarrow$ Sliced KSD $\downarrow$ log Z bias $\downarrow$
RDSMC 0.118 $\pm$ 0.002 0.013 $\pm$ 0.004 0.075 $\pm$ 0.005* 0.286 $\pm$ 0.068
RDSMC(Proposal) 0.091 $\pm$ 0.003 N/A 0.327 $\pm$ 0.033 N/A
RDMC 0.440 $\pm$ 0.003 N/A 0.129 $\pm$ 0.000 N/A
SLIPS 0.279 $\pm$ 0.003 N/A 0.076 $\pm$ 0.001 N/A
AIS 0.088 $\pm$ 0.003 0.004 $\pm$ 0.002* 0.087 $\pm$ 0.002 0.173 $\pm$ 0.010
SMC 0.088 $\pm$ 0.002* 0.006 $\pm$ 0.002 0.082 $\pm$ 0.004 0.169 $\pm$ 0.012*
SMS 0.282 $\pm$ 0.010 N/A 0.161 $\pm$ 0.000 N/A
Table 7: Ablation experiments: Rings and Funnel targets (mean $\pm$ standard error over 10 seeds). * indicates the best average results and bold indicates 95% confidence interval overlap with that of the best average result. AIS and SMC achieve the lowest average Radius TVD and $\log Z$ bias on Rings, and the lowest $\log Z$ bias on Funnel, with RDSMC remaining comparable when accounting for standard errors. RDSMC achieves the lowest Sliced KSD on Funnel.
We report the results in Table˜7. On synthetic targets, AIS and SMC achieve the lowest average Radius TVD and $\log Z$ bias on Rings and lowest average $\log Z$ bias on Funnel, while RDSMC achieves comparable results on these metrics and the lowest average Sliced KSD on Funnel.
Note that these observations differ slightly from those in the main experiments (Table˜1) due to changes in the hyperparameter settings. In particular, AIS and SMC no longer use target-variance scaling for initialization, which appears to improve their performance in this case.
In this new setting, AIS and SMC demonstrate more robust overall performance. While RDSMC remains competitive, it exhibits greater variability in the reported metrics (indicated by the larget standard errors). Nevertheless, RDSMC tends to outperform diffusion-based MC samplers that do not have SMC correction, including its Proposal variant, SLIPS, and RDMC.
The optimal hyperparameters for each method are selected based on the lowest Radius TVD for Rings and lowest Sliced KSD for Funnel, consistent with the procedure used in the main experiments.
Finally, we note that for RDSMC on the Rings target, when the hyperparameter $t_{\mathrm{start\text{-}resampling}}>0.8$, the magnitude of the estimated $\log Z$ becomes significantly larger than when $t_{\mathrm{start\text{-}resampling}}\leq 0.8$. For this reason, we tune $t_{\mathrm{start\text{-}resampling}}\in\{0,0.1,\ldots,0.8\}$, within which the estimated $\log Z$ values remain consistent.
Bayesian logistic regression.
Test LPPD $\uparrow$ Credit ($d=25$) Cancer ($d=31$) Ionosphere ($d=35$) Sonar ($d=61$)
RDSMC -94.54 $\pm$ 1.34 -10.59 $\pm$ 0.45 -25.97 $\pm$ 0.73 -18.54 $\pm$ 0.28
RDSMC(Proposal) -94.62 $\pm$ 0.06 -50.24 $\pm$ 0.16 -24.87 $\pm$ 0.02* -18.91 $\pm$ 0.01
RDMC -138.22 $\pm$ 0.35 -78.03 $\pm$ 0.12 -44.16 $\pm$ 0.07 -28.64 $\pm$ 0.03
SLIPS -92.42 $\pm$ 0.02* -10.41 $\pm$ 0.01 -25.22 $\pm$ 0.01 -18.40 $\pm$ 0.01*
AIS -92.91 $\pm$ 0.44 -10.13 $\pm$ 0.07 -25.09 $\pm$ 0.02 -18.41 $\pm$ 0.01
SMC -92.55 $\pm$ 0.05 -10.13 $\pm$ 0.01* -25.09 $\pm$ 0.02 -18.41 $\pm$ 0.01
SMS -98.00 $\pm$ 0.27 -20.47 $\pm$ 0.16 -26.17 $\pm$ 0.10 -23.29 $\pm$ 0.08
Table 8: Ablation experiments: Bayesian logistic regression with test log pointwise predictive density (Test LPPD, with mean $\pm$ standard error) averaged over 10 seeds. * indicates the best average result and bold indicates 95% confidence interval overlap with that of the best average result. RDSMC, AIS, SMC, and SLIPS achieve the best overall performance. RDSMC outperforms RDSMC (Proposal) and RDMC in most cases, highlighting the effectiveness of the SMC correction.
Different from the main experiments in Table˜8, we consider a different metric Test LPPD, the log pointwise predictive density on the test set, defined as follows
Test LPPD $=\sum_{(x,y)\in\mathcal{D}_{\textrm{test}}}\log\left(\int p(y_{i}\mid x_{i};w,b)\,p(w,b\mid\mathcal{D})\,\mathrm{d}w,b\right)$ (46)
where the likelihood $p(y_{i}\mid x_{i};w,b)$ is defined in Eq.˜43 and the posterior $p(w,b\mid\mathcal{D})$ is defined in Eq.˜44.
We use final (weighted) samples from each method to approximate the Test LPPD in Eq.˜46.
We report the results in Table˜8. We observe that RDSMC, AIS, SMC, and SLIPS achieve the best overall performance. Compared to SLIPS, RDSMC does not require target-variance-based initialization. Moreover, RDSMC outperforms RDSMC (Proposal) and RDMC in most cases, highlighting the effectiveness of the SMC correction. However, we observe that results for RDSMC exhibit large variance, as the case in the main experiments from Table˜2.
The optimal hyperparameters for each method are selected based on the highest LPPD on a heldout validation set, computed analogously to the Test LPPD.
E.5.3 Comparison to Particle Denoising Diffusion Sampler on the Funnel target
Finally, we present a comparison to the Particle Denoising Diffusion Sampler (PDDS) [phillips2024particle], which is also a diffusion-based SMC sampler. However, unlike RDSMC, PDDS requires training.
We evaluate PDDS on the Funnel target using 10 random seeds. Experiments are conducted on an NVIDIA RTX A6000 GPU, whereas our previous experiments use an NVIDIA A100 GPU. In addition, PDDS is implemented in JAX (following the official implementation at https://github.com/angusphillips/particle_denoising_diffusion_sampler# ), while RDSMC and the other baselines are implemented in PyTorch.
PDDS requires approximately $82.63\pm 0.43$ seconds for training and $15.08\pm 0.073$ seconds for final sampling of $4096$ particles, resulting in a total runtime of about $97.70\pm 0.47$ seconds (reported as mean $\pm$ standard error). The runtimes of our method and other training-free baselines are reported in Figure˜3, which are all below 25 seconds. However, due to differences in hardware and software frameworks, PDDS’s runtime is not directly comparable to those of the previous experiments.
PDDS achieves a $\log Z$ bias of $0.26\pm 0.04$ and a mean Sliced KSD of $0.10\pm 0.00$. Compared with the results in Table˜7, PDDS exhibits a slightly lower average $\log Z$ bias than RDSMC, but higher than AIS and SMC. Its Sliced KSD is also higher than that of RDSMC, AIS, and SMC.